### v2

# AICO: Feature Significance Tests for Supervised Learning

Abstract
The opacity of many supervised learning algorithms remains a key challenge, hindering scientific discovery and limiting broader deployment—particularly in high-stakes domains. This paper develops model- and distribution-agnostic significance tests to assess the influence of input features in any regression or classification algorithm. Our method evaluates a feature’s incremental contribution to model performance by masking its values across samples. Under the null hypothesis, the distribution of performance differences across a test set has a non-positive median. We construct a uniformly most powerful, randomized sign test for this median, yielding exact $p$-values for assessing feature significance and confidence intervals with exact coverage for estimating population-level feature importance. The approach requires minimal assumptions, avoids model retraining or auxiliary models, and remains computationally efficient even for large-scale, high-dimensional settings. Experiments on synthetic tasks validate its statistical and computational advantages, and applications to real-world data demonstrate its ability to inform high-stakes decisions with legal, economic, and regulatory implications.
1 Introduction
Supervised learning algorithms have become foundational tools across scientific and industrial domains, enabling breakthroughs and transforming key sectors. However, their opacity poses a major challenge, hindering scientific discovery and limiting adoption in high-stakes domains like finance, public policy and health care, where transparency is crucial.
This paper proposes novel testing methods for addressing the opacity of learning algorithms, focusing on the role of an algorithms’s input variables (features). Specifically, we develop a family of hypothesis tests to assess the significance of features in any supervised learning model, whether for regression or classification. The tests are model- and distribution-agnostic: they only require the ability to query the trained model—no assumptions about the model class, data distribution or model target (e.g. conditional mean, quantile) are needed. The computational efficiency of our tests is unprecedented. The procedure does not require retraining or surrogate models, making it well-suited for analyzing complex, high-dimensional algorithms that are costly to train.
The tests are based on a new approach for evaluating the impact of a feature. We quantify a feature’s incremental contribution to a user-defined score function, which measures a model’s predictive or decision-making performance on a test sample. Specifically, we compute the feature effect as the change in score when the model is given the actual value of the feature versus when the feature is masked with a baseline value. This masking approach, which we call Add-In Covariates (AICO), avoids the variance inflation (excess randomness) introduced by permutation methods and the high cost of retraining-based and surrogate-model-based strategies. The score function and masking baseline can be user-specified to align with specific interpretability goals and downstream model use cases; we discuss a number of choices and the associated tradeoffs.
The feature effect measures the impact of a feature’s variability on model score; it represents the importance of a feature at the level of an individual sample. A null-feature’s effect has a distribution across a collection of test samples whose median is non-positive. We construct a uniformly most powerful, randomized sign test for this median. Randomization at the boundary of the critical region ensures the test’s exactness, i.e., its size (type I error probability) matching the user-specified nominal level (e.g. $5\%$). The test delivers exact, randomized feature $p$-values that have, conditionally on the data, a uniform distribution on an interval which can be reported in practice. It is simple, robust and permits an efficient implementation. It requires minimal assumptions, ensuring its broad validity in practice. These properties allow users to make valid inferences regarding a model’s features without relying on large-sample approximations or parametric assumptions.
The median feature effect being tested is a natural measure of feature importance at the population level. We develop randomized, exact-coverage confidence intervals for the median that, together with the $p$-values, enable principled inference regarding feature influence. The availability of confidence intervals distinguishes the median from widely-used algorithmic measures of feature importance (see Section 1.1 below). Moreover, depending on the selection of the masking baseline, the median can have several useful properties. These include robustness to correlation between features, which is common in practice. Furthermore, for linear regression, the median is proportional to the square of the feature’s regression coefficient, which is the canonical measure of importance in the linear context. Finally, we provide conditions guaranteeing that a feature’s median being 0 is equivalent to it being conditionally independent of the response given the remaining features.
Extensive experiments on synthetic “known truth” tasks demonstrate the statistical and computational performance of our approach in nonlinear regression and classification settings. We compare our tests with the recently developed, model-agnostic and non-asymptotic tests of ?, ? and ?. We also compare with the model-X knockoff procedures of ? and ? for variable selection with false discovery control. Our test is found to be as statistically powerful as the alternatives, with minimal false discovery, while reducing computational costs by several orders of magnitude. Moreover, we show that our feature importance measure is robust to noise in the data, delivering stable feature importance rankings. This contrasts sharply with widely-used algorithmic feature importance measures including SHAP (?) which are found to be highly sensitive to noise in the data, leading to unstable feature rankings that make it difficult to draw reliable inferences regarding variable importance. Our findings complement the discussion of the theoretical caveats of algorithmic importance measures in ?, ? and others.
Empirical analyses demonstrate the ability of our framework to inform high-stakes decisions with legal, economic, and regulatory implications. In the context of a credit scoring task using large ensembles of gradient boosting machines, our approach identifies and ranks several significant features describing past payment behavior and social factors. Interestingly, we find that gender, marital status, and age of card holders (all of which are protected characteristics under US lending laws) are significant predictors of card default, after controlling for past payment behavior. In the context of a mortgage state transition prediction task using a recently proposed set-sequence model based on long-convolution architectures (?) and trained on extensive panel data, our test identifies time-varying macro-economic variables such as mortgage rates, as well as features summarizing past state transitions as significant predictors of borrower behavior.
The rest of this introduction discusses the related literature. Section 2 introduces the notion of feature effects while Section 3 develops the testing procedure. Section 4 analyzes the median feature effect as an importance measure dual to the test. Section 5 studies the performance of the test and importance measure in synthetic regression and classification settings. Section 6 deploys the test on two real-world data sets. Section 7 concludes. The proofs of technical results are given in an appendix. A Python package implementing our testing and importance procedures is available at https://github.com/ji-chartsiri/AICO.
1.1 Related literature
There is a large body of work developing algorithmic feature importance measures; see ? for an overview. Popular examples include measures based on Shapley values (e.g. SHAP, see ? and many others), measures based on local approximations (e.g. LIME, see ?), and measures obtained by permutation methods (e.g. ?, ?). These measures provide valuable insight into the role of features but unlike the feature importance measures delivered by our hypothesis testing approach, they typically do not offer statistical guarantees or $p$-values. We demonstrate that algorithmic measures can be highly susceptible to noise in the data so the lack of confidence intervals is concerning. Moreover, algorithmic measures such as SHAP can be biased in the presence of feature correlation (?). Our approach can deliver importance measures that are robust to feature correlation. Finally, we demonstrate that our method achieves orders-of-magnitude improvements in computational efficiency over widely used algorithmic importance approaches, such as SHAP and permutation scores.
The statistical and econometric literature has developed significance tests for the variables in nonparametric regression models. ?, ?, ?, ?, ? and others employ kernel methods to construct asymptotic-based testing procedures. ?, ? and ? and others construct gradient-based asymptotic tests using neural networks. Unlike these tests, ours are model-agnostic, non-asymptotic, deliver exact $p$-values and confidence intervals, and can accommodate a wide range of data structures.
Pursuing a model-agnostic approach, ?, ?, ?, ?, ?, ? and others develop non-asymptotic tests for a conditional independence hypothesis of feature significance in the model-X knockoff setting of ?. [^2] Conditional independence testing is challenging: ? show that no test of conditional independence can be uniformly valid against arbitrary alternatives. We show that, for an appropriate choice of the masking baseline, our null is generally weaker than the conditional independence null studied in these papers; equivalence holds under additional conditions. ? construct non-asymptotic tests based on the drop in model predictive accuracy due to removing features (LOCO: Leave-out-covariates). We, in contrast, test the improvement in model score due to masking a feature.
The computational requirements of the aforementioned model-agnostic testing procedures are a limiting factor. The LOCO test requires re-training the model for each feature to be tested, which can be prohibitive for feature-rich models that are costly to train. The model-X knockoff procedures require constructing an auxiliary model of the joint feature distribution, which can be as challenging and computationally costly as constructing the model to be tested. In fact, ? show that no algorithm can efficiently compute nontrivial knockoffs for arbitrary feature distributions. Our approach does not require any model re-training or the construction of auxiliary models, delivering an unprecedented degree of computational efficiency. Our numerical results demonstrate that the drastic gains in computational performance we achieve do not come at the expense of statistical performance.
2 Feature effect
We consider data $Z_{1},\ldots,Z_{m}$ with $Z_{i}=(X_{i},Y_{i})$. Here, $X_{i}=(X^{1}_{i},\ldots,X^{d}_{i})$ is a $d$-dimensional feature vector taking values in a set $\mathcal{X}$ and $Y_{i}$ is a response taking values in a set $\mathcal{Y}$. In a typical regression setting with tabular data, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$. We make no assumptions on the distribution of the $Z_{i}$’s.
We take as given a random split of $m]=\{1,\ldots,m\}$ into non-intersecting subsets $I_{1}$ and $I_{2}$, which are used for training/validation and testing purposes, respectively. This partitioning is standard in supervised learning settings. In our testing framework, data splitting facilitates post-selection inference.
We also take as given a model (algorithm) $f:\mathcal{X}\to\mathcal{M}$ trained on $\{Z_{i}\}_{i\in I_{1}}$, which is to be tested. We do not make any assumptions on the form of the modeling target; $f$ could represent $\mathbb{E}(Y_{i}|X_{i})$, a conditional quantile, or some other statistic of $Y_{i}$. We also do not impose assumptions on the training criterion (eg. $L_{2}$ error) or approach (eg. empirical risk minimization) being used for training.
We would like to test the statistical significance of a feature $\ell\in[d]$ under the model $f$. (Unless noted otherwise, $\ell$ is fixed.) In the special case that $f$ is linear, standard asymptotic procedures such as the $t$-test are available. We develop a simple, non-asymptotic and model-agnostic testing procedure that does not require a special model structure, and is valid much more broadly for models $f$ of essentially any form.
We consider the impact on model performance of a feature’s variability across the test set $I_{2}$. Performance is measured by a deterministic score function $S:\mathcal{M}\times\mathcal{Y}\to\mathbb{R}$, which relates model output to response. Specifically, for a given feature value $x\in\mathcal{X}$ and response value $y\in\mathcal{Y}$, the score $S(f(x),y)$ evaluates the quality of the model output $f(x)$ in view of $y$. Higher scores reflect better performance. A natural choice for $S$ is the negative of the loss function used for the training of $f$. Examples include the negative $L_{2}$ loss $S(m,y)=-(m-y)^{2}$ in a regression context, the negative pinball loss in a quantile regression context, or the negative (binary or multi-class) cross-entropy loss in a classification context. With these specifications $S$ measures predictive accuracy. Other specifications of $S$ can measure model performance (e.g. revenue) in the context of a downstream screening or decisioning application of $f$, for example the financial loss caused by a poor decision due to mis-classification. Finally, we can specify a function $S$ that does not depend on the second argument representing the response, to measure the behavior (e.g. sensitivity) of the model output on its own, without reference to the response.
We evaluate the effect on model score of a feature’s variability across the test set. For each $i\in I_{2}$ we construct from $X_{i}$ a baseline feature value $\underline{X}_{i}\in\mathcal{X}$ serving as a reference and a comparison feature value $\overline{X}_{i}\in\mathcal{X}$. In the baseline, the value of the feature to be tested is masked with a reference value whose choice is further discussed below. The comparison value $\overline{X}_{i}$ is unmasked so that the predicted response $f(\overline{X}_{i})$ can harness the actual sample $i$ value of the feature to be tested. We propose to measure the feature effect by the resulting score improvement.
Definition 2.1 (Feature Effect).
For a given score function $S$, baseline feature value $\underline{X}_{i}$ and comparison feature value $\overline{X}_{i}$, the feature effect on test sample $i\in I_{2}$ is the score delta
$\Delta_{i}=S(f(\overline{X}_{i}),Y_{i})-S(f(\underline{X}_{i}),Y_{i}).$
Note that masking feature values for purposes of measuring a feature effect has significant advantages over alternative approaches such as randomly shuffling feature values (e.g. ? and ?) or dropping a feature from the feature vector. Shuffling introduces excess variance into the measurement procedure and dropping features requires training sub-models, which (drastically) reduces the computational efficiency of the measurement procedure. Our masking approach avoids these issues.
2.1 Baseline and comparison features
Along with the selection of the score function $S$, the selection of baseline and comparison feature values enables users to tailor the feature effect to the model and data structures at hand as well as the testing objectives. We provide a conceptional framework and several examples. Throughout, fix a feature $\ell\in[d]$ that is to be tested; our notation usually suppresses the dependence on $\ell$. We start by selecting a non-empty set $\underline{I}\subseteq I_{1}$ of training samples that will be used to construct the baseline $\underline{X}_{i}$ for the test sample $i\in I_{2}$. In a second step, we aggregate the samples in $\underline{I}$ to obtain $\underline{X}_{i}$. Alternative aggregators include the mean, a quantile or a mode of $\{X_{k}:k\in\underline{I}\}$.
Example 2.1.
Take $\underline{I}=I_{1}$ and let $\underline{X}_{i}=(v^{1},\ldots,v^{d})$ for all $i\in I_{2}$ where $v^{j}$ is the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ for each $j\in[d]$. The predicted response $f(\underline{X}_{i})$ is constant across the test set. If feature $\ell=1$ is to be tested, select the comparison value $\overline{X}_{i}=(X_{i}^{1},v^{2},\ldots,v^{d})$. The predicted response $f(\overline{X}_{i})$ harnesses feature 1’s actual sample value, $X_{i}^{1}$. The feature effect $\Delta_{i}$ represents the associated score improvement.
An alternative choice is $\underline{X}_{i}=(v^{1},X^{2}_{i},\ldots,X^{d}_{i})$ and $\overline{X}_{i}=X_{i}=(X_{i}^{1},\ldots,X^{d}_{i})$, again supposing feature $\ell=1$ is to be tested. The predicted baseline response $f(\underline{X}_{i})$ is no longer constant across the test set, and accounts for the variation of features $X^{2}_{i},\ldots,X^{d}_{i}$. Here $\Delta_{i}$ measures a conditional feature effect.
Choices other than the mean may be appropriate for $v^{j}$. For example, in the case of a one-hot feature $X^{j}_{i}\in\{0,1\}$ which might encode a categorical variable, sensible choices for $v^{j}$ are a mode of $\{X^{j}_{k}:k\in\underline{I}\}$ or the value $1-X^{j}_{i}$. [^3] See ? for extensions.
Consider choices of $\underline{I}$ beyond the entire training set $I_{1}$. We can pick specific samples from $I_{1}$ by first fixing a reference point $r$ in $\mathcal{Y}$, $\mathcal{M}$, or $\mathcal{X}$ and then selecting the samples from $I_{1}$ that lie in or outside a specified neighborhood of $r$. The point $r$ can represent the (weighted) mean, mode or a $\lambda$-quantile of the responses $\{Y_{k}:k\in I_{1}\}$, the predicted responses $\{f(X_{k}):k\in I_{1}\}$, or the values of the feature $\ell$ to be tested, $\{X^{\ell}_{k}:k\in I_{1}\}$. Potentially interesting special cases are quantiles for $\lambda\in\{0,1\}$. Concretely, let $Y_{r}\in\mathcal{Y}$ be a reference response value and $d$ be a distance measure on $\mathcal{Y}$. We can now take $\underline{I}$ as the set of samples $k\in I_{1}$ for which $d(Y_{k},Y_{r})$ does not exceed or surpasses some threshold value. Other criteria can be envisioned, for example samples $k\in I_{1}$ for which $Y_{k}$ exceeds or falls short of $Y_{r}$. A similar approach applies to reference points in $\mathcal{M}$ or $\mathcal{X}$.
Example 2.2.
Consider the reference point $\min_{k\in I_{1}}f(Y_{k})$ in $\mathcal{M}$, which is a quantile of the predicted responses $\{f(Y_{k}):k\in I_{1}\}$ at level $\lambda=0$. For example, in a binary soft classification setting such as fraud detection or credit scoring where $Y_{i}=0$ represents the preferred (beneficial) outcome, this reference point represents the training samples with minimum predicted risk. [^4] See ?. We can take $\underline{I}=\{\mbox{argmin}_{k\in I_{1}}f(Y_{k})\}$ as the set of these samples and define $\underline{X}_{i}$ and $\overline{X}_{i}$ as in Example 2.1. The associated feature effect $\Delta_{i}$ represents the score delta relative to the minimum risk baseline. It measures how much excess risk a feature explains.
A few basic principles guide the selection of reference points and the construction of the set $\underline{I}\subseteq I_{1}$. We wish to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, to avoid feature effects $\Delta_{i}$ that are biased low which result in significance tests with low power.555? discuss these issues in the context of image classification. Similarly, we want to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$. Finally, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space. It is important to select an appropriate reference point and rule used to aggregate across $\underline{I}$ (see Example 2.1). Alternatively, we might select an appropriate baseline value from the set $\{X_{k}:k\in I_{1}\}$, namely an element that is near the raw off-manifold value emerging from the baseline selection procedure.
For clarity, the discussion above has focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$. The approach naturally extends to testing multiple features, i.e. subsets of $[d]$, which facilitates the exploration of higher-order effects including feature interactions. We also allow for the selection and aggregation of alternative baselines for a given sample. A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
2.2 Panel data
Panel data settings, where a collection of units are observed over some time period, are prevalent in many domains. While our approach assumes independent samples, a simple extension—grouping—allow us to treat some dependent settings like panel data. The idea is to aggregate feature effects within each unit of dependence, so that dependence is contained within groups with inter-group dependencies being mitigated.
Concretely, consider panel data $\{Z_{i,t}\}$, where $i=1,\dots,m$ indexes units and $t=1,\dots,T$ indexes time periods. We take as given a unit-level random split of the index set $m]=\{1,\dots,m\}$ into disjoint training and testing subsets $I_{1}$ and $I_{2}$, respectively, such that each unit, along with its entire time series, is assigned exclusively to either $I_{1}$ or $I_{2}$. We treat the test set $I_{2}$ as a collection of $|I_{2}|T$ samples and compute the per-unit-time feature effects $\Delta_{i,t}$ for each $(i,t)\in I_{2}$. We then aggregate these effects within each unit: $\Delta_{i}=\frac{1}{T}\sum_{t=1}^{T}\Delta_{i,t}$.
For purposes of computing baseline and comparison features in this group setting, we also treat the training data $I_{1}$ at a unit level. For each $(i,t)\in I_{1}$ and feature $j\in[d]$, the per-unit-time samples are aggregated to obtain grouped feature samples $X_{i}^{j}=a_{i}^{j}(X^{j}_{i,1},\dots,X^{j}_{i,T})$, where $a_{i}^{j}$ is an aggregation function for the time series associated with unit $i$ (e.g. the mean, median, mode).
The grouping approach naturally extends to unbalanced panel data, where subjects may be observed at different time points or for different durations. In such cases, the average for each subject is computed only over the observed periods.
More broadly, the grouping approach can be adapted to other dependence structures beyond temporal ones. For instance, for spatially dependent data, one could group feature effects by region or neighborhood, under the assumption that neighborhoods are (approximately) independent. Similarly, in settings with repeated measurements, grouping by the underlying entity or context is effective.
3 Hypothesis testing
To construct a hypothesis test, we consider the empirical distribution of the feature effect $\Delta_{i}$ across the test samples $i\in I_{2}$ (the feature $\ell\in[d]$ to be tested remains fixed). If that distribution has a non-positive center, then the feature is deemed insignificant. A large positive center is evidence of a feature being significant. There is a battery of alternative tests for the mean, median and other statistics, including $t$, sign, Wicoxon, and permutation tests. We propose a simple sign test, which requires minimal assumptions on the data yet offers meaningful asymptotic relative efficiency vs. a $t$-test.
Assumption 3.1.
The feature effects $\{\Delta_{i}:i\in I_{2}\}$ are independent draws from a conditional distribution $\mathbb{P}_{M}$ given the training data $I_{1}$ with unique median $M\in\mathbb{R}$. The associated distribution function $F_{\Delta}$ is continuous. [^6] A sufficient condition for the uniqueness of $M$ is strict monotonicity of $F_{\Delta}$ in a neighborhood of $M$.
Let $B_{n,p}$ be a Binomial-$(n,p)$ variable and let $q_{\lambda}(B_{n,p})=\inf\{x\geq 0:P(B_{n,p}\leq x)\geq\lambda\}$ be its (lower) quantile at $\lambda\in(0,1)$. Let $N=|I_{2}|$ be the test sample size. Moreover, let $\Pi_{N}(n)=\sum_{i=0}^{n}\pi_{N}(i)$ be the CDF of $B_{N,1/2}$, where $\pi_{N}(i)={N\choose i}2^{-N}$ for $i=0,1,\ldots,N$.
Proposition 3.2.
Under Assumption 3.1, consider the hypotheses
$H_{0}:M\leq M_{0}\quad vs.\quad H_{1}:M>M_{0}$ (1)
for some $M_{0}\in\mathbb{R}$. A Uniformly Most Powerful test of size $\alpha\in(0,1)$ rejects $H_{0}$ with probability $\phi_{N}(n_{+}(M_{0}),\alpha)$ where $n_{+}(x)=\#\{i\in I_{2}:\Delta_{i}>x\}$ is the number of feature effect samples in the test set exceeding $x\in\mathbb{R}$ and
$\phi_{N}(n,\alpha)=1\{n>T_{N,\alpha}\}+\gamma_{N}(n,\alpha)1\{n=T_{N,\alpha}\}$ (2)
where $T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ and
$\gamma_{N}(n,\alpha)=\frac{\Pi_{N}(n)-(1-\alpha)}{\pi_{N}(n)},\qquad n=0,1,\ldots,N.$ (3)
Input:
-0.5em
  • Model $f$ trained on $\{X_{i},Y_{i}\}_{i\in I_{1}}$
  • Baseline feature values $\{\underline{X}_{i}\}_{i\in I_{2}}$ for $\ell$
  • Comparison feature values $\{\overline{X}_{i}\}_{i\in I_{2}}$ for $\ell$
  • Response values $\{Y_{i}\}_{i\in I_{2}}$
  • Significance level $\alpha\in(0,1)$
  • Score function $S$
-0.5em Output:
-0.5em
  • Test conclusion (True = Significant at level $\alpha$, False = Insignificant at level $\alpha$)
-0.5em
Compute $\Delta_{i}=S(f(\overline{X}_{i}),Y_{i})-S(f(\underline{X}_{i}),Y_{i})$ for all $i\in I_{2}$ ;
Compute the test statistic $n_{+}=\#\{i\in I_{2}:\Delta_{i}>M_{0}\}$ ;
Compute the threshold $T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ ;
if $n_{+}>T_{N,\alpha}$ then
 return True;
else if $n_{+}=T_{N,\alpha}$ then
   Generate $U\sim U(0,1)$ ;
   Compute the probability $\gamma_{N}(n_{+},\alpha)$ using (3) ;
 if $U\leq\gamma_{N}(n_{+},\alpha)$ then
    return True ;
return False
0.5em
Algorithm 1 AICO Test algorithm for feature $\ell\in[d]$.
For brevity, we often suppress the dependence of the test statistic $n_{+}$ on $M_{0}$. The test rejects $H_{0}$ with probability one if $n_{+}$, the number of feature effects in the test set larger than $M_{0}$, exceeds the $(1-\alpha)$-quantile of the Binomial-$(N,1/2)$ distribution. The test randomizes in case $n_{+}$ hits the boundary of the rejection region. Relative to a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$, randomization renders the test more aggressive; it ensures maximum power and guarantees exactness in the sense that
$\mathbb{E}_{M_{0}}(\phi(n_{+},\alpha))=\alpha,\quad M_{0}\in\mathbb{R}.$ (4)
Algorithm 1 details the implementation of the test. Our applications of the test in Sections 5 and 6 below will use the value $M_{0}=0$ in (1).
In the context of our randomized test the notion of $p$-value is ambiguous. A conventional $p$-value could be defined as $\inf\{\alpha:\phi_{N}(n_{+},\alpha)=1\}$, the smallest level of significance $\alpha$ at which $H_{0}$ is rejected with probability one, which is $1-\Pi_{N}(n_{+}-1)$. (This is also a $p$-value for a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$.) However, this formulation ignores the randomness of the UMP test’s decision rule. We follow ? and consider a randomized $p$-value $p(n_{+})$ with conditional distribution function given by
$\mathbb{P}_{M_{0}}(p(n_{+})\leq\alpha\,|\,n_{+})=\phi_{N}(n_{+},\alpha),\quad\alpha\in(0,1).$
From (2), this function represents a uniform distribution on an interval $(a_{N}(n_{+}),b_{N}(n_{+}))$ of length $\pi_{N}(n_{+})$ where
$a_{N}(n)$ $=P(B_{N,1/2}>n)=1-\Pi_{N}(n),$
$b_{N}(n)$ $=P(B_{N,1/2}\geq n)=1-\Pi_{N}(n-1),$
the right endpoint being the conventional $p$-value. By iterated expectations and (4), we have
$\mathbb{P}_{M_{0}}(p(n_{+})\leq\alpha)=\mathbb{E}_{M_{0}}(\phi_{N}(n_{+},\alpha))=\alpha,\quad M_{0}\in\mathbb{R},$
meaning that $p(n_{+})\sim U(0,1)$. In this sense, the randomized $p$-value $p(n_{+})$ is exact, mirroring the exactness of the test’s decision rule (2). If $p(n_{+})$ is concentrated below a given level $\alpha$, or equivalently, if the conventional $p$-value $b_{N}(n_{+})\leq\alpha$, we reject the null. If $p(n_{+})$ is concentrated above $\alpha$, or equivalently, if $\alpha<a_{N}(n_{+})$, we retain the null. Else the evidence is equivocal; we thus randomize by drawing a sample $u$ from $U(a_{N}(n_{+}),b_{N}(n_{+}))$ and reject if $u\leq\alpha$. See Figure 1 for a visualization. In practice, it suffices to report the interval $(a_{N}(n_{+}),b_{N}(n_{+}))$ to fully characterize the $p$-value; there is often no need to force a decision.
[…]
Figure 1: Distribution of randomized $p$-value and decision intervals.
The power function $H_{N,\alpha}(M)$ of the test is given by
$H_{N,\alpha}(M)$ $=\mathbb{E}_{M}(\phi_{N}(n_{+}(M_{0}),\alpha))$
$=\mathbb{P}_{M}(n_{+}(M_{0})>T_{N,\alpha})+\gamma_{N}(T_{N,\alpha},\alpha)\mathbb{P}_{M}(n_{+}(M_{0})=T_{N,\alpha})$
$=P(B_{N,p}>T_{N,\alpha})+\gamma_{N}(T_{N,\alpha},\alpha)P(B_{N,p}=T_{N,\alpha})$ (5)
for values $M\in\mathbb{R}$, where $p=\mathbb{P}_{M}(\Delta_{i}>M_{0})$. The binomial probabilities in (3) are explicit. Since the test is uniformly most powerful, the function $H_{N,\alpha}$ is increasing for fixed $N$ and $\alpha$. We can invert that function to determine the size of the test set $N=|I_{2}|$ required to achieve a desired level of power. The parameter $p$ can be estimated using a pilot sample or other approach, as discussed by ?. Figure 2 plots the test sample size as a function of test power for several alternative values of $p$.
Figure 2: Test sample size NN vs. power $H_{N,\alpha}$ for $\alpha=5\%$ and several values of $p>1/2$.
The multiple testing problem can also be addressed in the randomized setting. A standard Bonferroni adjustment with test function $\phi_{N}(n,\alpha/d)$ and $p$-value being conditionally uniform on the interval $(da_{N}(n_{+})\wedge 1,db_{N}(n_{+})\wedge 1)$, where $d$ is the dimension of the feature vector, controls the family-wise error rate (FWER, the probability of having at least one false discovery), see ?. For a less conservative approach, one can also control the False Discovery Rate (?). ? develop a formula and an algorithm for the adjusted rejection probability. ? proposes an alternative Monte Carlo approach. ? provide a comparison of methods, finding that the Monte Carlo approach outperforms its alternatives in terms of test power.
4 Measuring feature importance
We require measures of feature importance that complement the feature $p$-values delivered by our test. A local measure of feature importance is given by the feature effect $\Delta_{i}$, which represents the impact of a feature at the sample level. The median $M$ of the (conditional) feature effect distribution $F_{\Delta}$ is a natural measure of feature importance at the population level. Under Assumption 3.1, a consistent estimator of $M$ is the empirical median
$\widehat{M}=\mbox{Median}\{\Delta_{i}:i\in I_{2}\},$ (6)
which satisfies $\mathbb{P}(\widehat{M}<M)=\mathbb{P}(\widehat{M}>M)$ for odd $N=|I_{2}|$ and approximate equality holding for even $N$. Thus (6) is (approximately) as likely to overestimate $M$ as to underestimate it. Moreover, the distribution of $\widehat{M}$ is continuous. The estimator $\widehat{M}$ is unbiased if $F_{\Delta}$ is symmetric about $M$; see Section 4.4 of ?.
We construct one-sided, randomized confidence intervals (CIs) for $M$ by inverting the test (2). Let $\{\Delta_{(i)}\}$ denote the ordered sequence of $\{\Delta_{i}\}$; we take $\Delta_{(0)}=$-$\infty$ and $\Delta_{(N+1)}=\infty$. Let $\mathbb{P}_{M}$ denote the sampling distribution of the $\{\Delta_{i}\}$ when $M$ is the true median.
Proposition 4.1.
Let $U$ be an independent $U(0,1)$ variable. Define a randomized CI by
$C_{N}(\alpha;U)=\begin{cases}[\Delta_{(N-T_{N,\alpha})},\infty),&1-\gamma_{N}(T_{N,\alpha},\alpha)\geq U\\ [\Delta_{(N-T_{N,\alpha}+1)},\infty),&\text{else}\end{cases}$ (7)
where $T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ is the boundary of the test’s rejection region and $\gamma_{N}(n,\alpha)$ is the rejection probability (3). Under Assumption 3.1, the CI (7) has exact coverage:
$\mathbb{P}_{M}(M\in C_{N}(\alpha;U))=1-\alpha,\quad\alpha\in(0,1).$ (8)
In analogy to the randomized $p$-value, the randomized CI (7) inherits the exactness of the test’s decision rule (2). If $0<\Delta_{(N-T_{N,\alpha})}$ we reject the null for $M_{0}=0$. If $\Delta_{(N-T_{N,\alpha}+1)}\leq 0$ we retain the null. Else we reject with probability $\gamma_{N}(T_{N,\alpha},\alpha)$ (that is, draw a sample $u$ from $U(0,1)$ and reject if $u\leq\gamma_{N}(T_{N,\alpha},\alpha)$). [^7] For comparison, a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$ has CI $\Delta_{(N-T_{N,\alpha})},\infty)$ and rejects only if $M_{0}=0<\Delta_{(N-T_{N,\alpha})}$, implying a more conservative approach. In practice, it suffices to report the interval endpoints along with the rejection probability.
[…]
Figure 3: Rejection probability vs. CI endpoints.
Standard arguments lead to alternative two-sided CIs, which are not, however, dual to the test (2). Observe that $\mathbb{P}_{M}(\Delta_{(i)}\leq M\leq\Delta_{(i+1)})=\pi_{N}(i)$ for $i=0,1,\ldots,N$. Using these probabilities, we can combine intervals $(\Delta_{(i)},\Delta_{(i+1)})$ so as to attain a desired confidence coefficient. Choosing the intervals symmetrically, we obtain intervals
$c_{m,N}=[\Delta_{(1+m)},\Delta_{(N-m)}]$ (9)
satisfying $\mathbb{P}_{M}(M\in c_{m,N})=1-2\Pi_{N}(m)$, where we select $m\in\mathbb{N}$ so that $\Pi_{N}(m)\leq\alpha/2$ for given $\alpha\in(0,1)$. Then, $\mathbb{P}_{M}(M\in c_{m,N})\geq 1-\alpha$, implying a conservative approach. ? extend this argument using randomization, resulting in CIs with exact $(1-\alpha)$ coverage and minimum expected length. Alternative asymptotic CIs can be obtained by bootstrapping (e.g. ? and ?), empirical likelihood (?), sectioning (?) and other methods.
We develop several useful properties of the population-level feature importance measure defined by the median $M$. To this end, let $Z=(X,Y)\sim\mathbb{P}$ where $\mathbb{P}$ denotes the population law from which the observed data points $Z_{i}$ are independently drawn. For concreteness, consider a model for estimating the conditional mean, and take $\mu(X)=\mathbb{E}(Y|X)$ and the score function $S$ as given. The (population) feature effect is given by $\Delta=S(\mu(\overline{X}),Y)-S(\mu(\underline{X}),Y)$ where $\underline{X}$ and $\overline{X}$ are suitable baseline and comparison features, respectively.
As discussed in Section 2.1, the selection of $\underline{X}$ and $\overline{X}$ offers some degrees of freedom. Consider the basic choice from Example 2.1,
$\underline{X}=\mathbb{E}(X)\quad\text{and}\quad\overline{X}=(X^{1},\mathbb{E}(X^{2}),\ldots,\mathbb{E}(X^{d})),$ (10)
where for convenience we assume, without loss of generality, feature $X^{1}$ is to be tested. First note that for the choice (10) and given $\mu$ and $S$, the feature effect $\Delta$ is completely determined by the marginal laws of the features and the joint law of the response $Y$ and $X^{1}$. Thus the feature dependence structure does not matter; the importance measure $M$ generated by (10) is robust to feature correlation. This is a remarkable property given the sensitivity of popular importance measures including SHAP and LOCO to correlation among features (e.g. ?). The correlation sensitivity induces a bias in these measures, which can lead to dubious importance rankings.
Next we explore the relationship between $M$ and conditional independence. ?, ?, ?, ?, ? and others develop tests for a conditional independence hypothesis of feature significance in the model-X knockoff setting of ?. We show that the hypothesis we test ($M\leq M_{0}=0$) is generally weaker than an alternative conditional independence hypothesis. Below, we use the notation $X^{-1}=(X^{2},\ldots,X^{d})$.
Proposition 4.2.
[…]
$Y=\mu(X)+\epsilon,\;\text{where}\;\mu(X)=\beta\cdot X\;\text{and}\;\beta=(\beta^{1},...,\beta^{d})\in\mathbb{R}^{d}.$ (11)
Before stating a converse of Proposition 4.2 under (11), we note the following “linear agreement” property: under (11) (and symmetry of the distribution of $\Delta$), the median $M$ measuring feature importance is proportional to the square of the regression coefficient, which is the canonical measure of feature importance in a linear model. The proportionality factor is given by the variance of the feature.
Proposition 4.3.
Consider the linear model (11). Select (10) and a standard score function $S(m,y)=-(m-y)^{2}$. If $\mathbb{E}(X^{-1}|X^{1})=\mathbb{E}(X^{-1})$ and $\mathbb{E}(\epsilon|X^{1})=0$, we have
$\mathbb{E}(\Delta)=(\beta^{1})^{2}var(X^{1}).$
Proposition $4.$
[…]
5 Synthetic tasks
This section illustrates the robust performance of our test and feature importance measure on synthetic “known truth” regression and classification tasks using neural networks. It also benchmarks performance against existing model-agnostic tests and variable selection procedures, as well as alternative model-specific and -agnostic feature importance measures.
5.1 Regression
We consider a feature vector $X=(X^{1},\ldots,X^{d})\subseteq\mathbb{R}^{d}$ of dimension $d=19$ and a response $Y\in\mathbb{R}$ of the form
$Y=\mu(X)+\epsilon,\quad\epsilon\sim N(0,1),$ (12)
where the error $\epsilon$ is independent of $X$ and the conditional mean function $\mu$ is given by
$\mu(x)=3$ $+4x_{1}+x_{1}x_{2}+3x_{3}^{2}+2x_{4}x_{5}+6x_{6}+2\sin(x_{7})$
$+\exp(x_{8})+5x_{9}+3x_{10}+4x_{11}+5x_{12}.$ (13)
The features $X_{1}$ and $X_{6}$ are jointly normal with mean zero, unit variances and covariances equal to $0.85$. The features $X_{2},\ldots,X_{5},X_{7}$ are independent $N(0,1)$, $X_{8}$ is uniform on [1,1][-1,1], $X_{9}=1\{X_{2}+Z<0\}$ where $Z\sim N(0,1)$, $X_{10}$ is Poisson with mean 3 and $X_{11}$ and $X_{12}$ are independent variables with a $t$-distribution and 5 degrees of freedom. The non-influential variables $X_{13}$ and $X_{14}$ are independent $N(0,1)$, $(X_{15},X_{16})$ are jointly normal with mean zero, unit variances and covariances equal to $0.85$, and $X_{17},X_{18}$ and $X_{19}$ are independent variables with a $t$-distribution and 5 degrees of freedom.
We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. We train a single-layer neural network model $f$ of the conditional mean $\mu(X)=\mathbb{E}(Y|X)$. Specifically,
$f(x)=b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x),$ (14)
where $K$ is the number of units, $\psi(y)=1/(1+e^{-y})$ is the activation function, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters to be estimated. The training procedure uses a standard $L_{2}$ loss criterion, $L_{2}$ regularization (of the weights $b_{k}$), the Adam optimization method, a batch size of 32, and 500 epochs with early stopping. We iterate across five random initializations. A validation set ($25\%$ of $I_{1}$, not used for training) and a Bayesian approach are used to optimize the regularization weight, the learning rate and the number of units, $K$. The optimal weight is $7\cdot 10^{-4}$, the optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$. The $R^{2}$ on the training (and test) data exceeds $99\%$.
The testing procedure uses a score function $S(m,y)=-(m-y)^{2}$ consistent with the $L_{2}$ training criterion. We use the basic baseline and comparison feature specifications discussed in Example 2.1. That is, for testing feature $\ell\in[d]$ we take
$\underline{X}_{i}$ $=(X^{1}_{i},\ldots,X^{\ell-1}_{i},v^{\ell}_{i},X^{\ell+1}_{i},\ldots,X^{d}_{i})$
$\overline{X}_{i}$ $=X_{i}=(X_{i}^{1},\ldots,X^{d}_{i})$ (15)
where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ in case feature $\ell$ is continuous. For the one-hot feature $\ell=9$, we use $v^{\ell}_{i}=1-X^{\ell}_{i}$ and for the Poisson feature $\ell=10$, we use the “alternative mode” $v^{\ell}_{i}={\arg\max}_{x\in\mathbb{N}\setminus\{X^{\ell}_{i}\}}p^{\ell}(x|I_{1})$ where $p^{\ell}(\cdot|I_{1})$ is the probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$. The latter choice avoids baseline and comparison features being identical for too many samples, which is an issue arising with the standard mode. Our specifications yield baseline values in the support of $X^{9}$ and $X^{10}$.
$\ell$ AICO CPT HRT LOCO Deep Gaussian
Knockoff Knockoff
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 10 10 10 10 10 10 10 10
2 10 10 10 10 10 10 10 10 10 10 10 10
3 10 10 10 10 10 10 10 10 10 10 10 10
4 10 10 10 10 10 10 10 10 8 8 10 10
5 10 10 10 10 10 10 10 10 8 8 10 10
6 10 10 10 10 10 10 10 10 10 10 10 10
7 10 10 10 10 10 10 10 10 10 10 10 10
8 10 10 10 10 10 10 10 10 9 9 10 10
9 10 10 10 10 10 10 10 10 10 10 10 10
10 10 10 10 10 10 10 10 10 10 10 10 10
11 10 10 10 10 10 10 10 10 10 10 10 10
12 10 10 10 10 10 10 10 10 10 10 10 10
13 0 0 3 2 0 0 0 0 3 3 4 4
14 0 0 1 1 1 1 1 1 2 2 3 3
15 0 0 2 2 0 0 0 0 0 0 4 4
16 0 0 4 4 1 1 0 0 1 1 3 3
17 0 0 0 0 1 1 1 0 2 2 3 3
18 0 0 0 0 1 1 0 0 0 0 2 2
19 0 0 1 0 0 0 0 0 0 0 2 2
Table 1: Regression model test results for 10 independent training/testing experiments at two standard confidence levels ($\alpha=5\%$ and $\alpha=1\%$). The table compares AICO, CPT, HRT, LOCO, Deep Knockoff, and Gaussian Knockoff procedures, showing the total number of rejections across 10 trials. The test set size is $N=5\cdot 10^{5}$.
We run 10 independent training/testing experiments to assess the behavior of the test. The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.888Limited numerical precision can lead to ties, i.e. samples $\Delta_{i}$ whose true values are distinct are being rounded to the same value. If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$. Another numerical artifact is samples $\Delta_{i}$ with true, non-zero value being rounded to 0. We treat these samples as if they indeed had non-positive values not contributing to $n_{+}$, implying a conservative approach. Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them. See also ? for another approach. The test has perfect power and does not produce false discoveries. Feature correlation (e.g. $\ell=1,6$) and discrete feature distributions ($\ell=9,10$) do not affect performance. The test also correctly detects the significance of features $\ell=2,4,5$ which influence the response only through product terms. Additional experiments show robustness to alternative baseline specifications (e.g. taking $v^{\ell}_{i}$ as the training mean also for the discrete features $\ell=9,10$, and using the median instead of the mean to construct $v^{\ell}_{i}$). We also verify robustness with respect to the score function $S$ including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
[…]
Table 1 reports the results of the aforementioned alternative procedures for 10 independent training/testing experiments. CPT, HRT and LOCO deliver the same power as AICO but come with an increased false discovery rate (some but not all of these false discoveries can be addressed by a standard Bonferroni correction). The Gaussian Knockoff procedure selects the significant features 1 through 12 in each independent experiment. However, it does generate some false discoveries (which is consistent with control of the expected false discovery rate). The Deep Knockoff procedure reduces false discoveries at the expense of lower power. This approach has slight difficulty identifying the features entering the model (5.1) through the product term $2x_{4}x_{5}$.
We consider the performance of our AICO test as a function of the size $N=|I_{2}|$ of the test set. Table 2 reports the results for $N\in\{10^{5},10^{4},10^{3},500,100\}$; the case $N=5\cdot 10^{5}$ is reported in Table 1. The test offers maximum power even for $N$ as small as 500. The power deteriorates moderately for smaller $N$; false discoveries rise only very slowly. This validates the performance of the test even for small samples.
AICO
$\ell$ $N=10^{5}$ $N=10^{4}$ $N=10^{3}$ $N=500$ $N=250$ $N=100$
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 10 10 10 10 10 10 10 10
2 10 10 10 10 10 10 10 10 10 9 8 5
3 10 10 10 10 10 10 10 10 10 10 10 10
4 10 10 10 10 10 10 10 10 10 10 9 9
5 10 10 10 10 10 10 10 10 10 10 10 9
6 10 10 10 10 10 10 10 10 10 10 10 10
7 10 10 10 10 10 10 10 10 10 10 10 10
8 10 10 10 10 10 10 10 10 9 8 8 7
9 10 10 10 10 10 10 10 10 10 10 10 10
10 10 10 10 10 10 10 10 10 10 10 10 10
11 10 10 10 10 10 10 10 10 10 10 10 10
12 10 10 10 10 10 10 10 10 10 10 10 10
13 0 0 0 0 0 0 0 0 2 1 0 0
14 0 0 0 0 0 0 1 0 0 0 1 0
15 0 0 1 0 1 1 0 0 1 0 1 0
16 0 0 0 0 0 0 1 1 0 0 1 0
17 1 0 0 0 1 0 0 0 0 0 0 0
18 0 0 1 0 0 0 0 0 0 0 0 0
19 0 0 1 0 0 0 0 0 0 0 2 2
Table 2: Regression model test results for 10 independent training/testing experiments, evaluated at two confidence levels ($\alpha=5\%$ and $1\%$) for several test sample sizes $N$. We report the number of times the null was rejected out of 10 trials for each covariate $\ell\in[d]$.
We turn to measuring feature importance. Table 3 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen. It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities. Table 3 also provides the distribution of the randomized $p$-value. [^10] Ties of samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing CIs and coverage probabilities. This is consistent with their treatment in the testing procedure (footnote 8).
AICO, Level $\alpha=1\%$
$p$-value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
Rank Feature Median P(Reject) Lower Upper 98.8% 1.2% Coverage Lower Upper Coverage
1 $X_{9}$ 24.993 1.000 0.000 0.000 24.951 24.951 0.99 24.947 25.040 0.99007
2 $X_{6}$ 15.788 1.000 0.000 0.000 15.661 15.661 0.99 15.647 15.938 0.99007
3 $X_{10}$ 14.572 1.000 0.000 0.000 14.511 14.511 0.99 14.506 14.640 0.99007
4 $X_{12}$ 12.898 1.000 0.000 0.000 12.789 12.789 0.99 12.775 13.022 0.99007
5 $X_{11}$ 8.137 1.000 0.000 0.000 8.063 8.063 0.99 8.054 8.218 0.99007
6 $X_{1}$ 6.210 1.000 0.000 0.000 6.155 6.155 0.99 6.148 6.273 0.99007
7 $X_{3}$ 1.697 1.000 0.000 0.000 1.669 1.669 0.99 1.666 1.730 0.99007
8 $X_{7}$ 0.953 1.000 0.000 0.000 0.941 0.941 0.99 0.940 0.964 0.99007
9 $X_{4}$ 0.416 1.000 0.000 0.000 0.409 0.409 0.99 0.409 0.424 0.99007
10 $X_{5}$ 0.416 1.000 0.000 0.000 0.409 0.409 0.99 0.408 0.424 0.99007
11 $X_{8}$ 0.143 1.000 0.000 0.000 0.140 0.140 0.99 0.139 0.146 0.99007
12 $X_{2}$ 0.084 1.000 0.000 0.000 0.082 0.082 0.99 0.082 0.086 0.99007
- $X_{13}$ -1.154e-07 0.000 0.880 0.881 -2.442e-06 -2.437e-06 0.99 -2.708e-06 1.478e-06 0.99007
- $X_{14}$ -2.054e-07 0.000 0.928 0.928 -2.176e-06 -2.175e-06 0.99 -2.389e-06 9.014e-07 0.99007
- $X_{15}$ -3.060e-07 0.000 0.845 0.846 -3.339e-06 -3.338e-06 0.99 -3.639e-06 2.127e-06 0.99007
- $X_{16}$ -5.512e-07 0.000 0.937 0.937 -2.966e-06 -2.965e-06 0.99 -3.275e-06 9.626e-07 0.99007
- $X_{17}$ -1.340e-06 0.000 0.996 0.997 -3.773e-06 -3.773e-06 0.99 -4.022e-06 0.000 0.99150
- $X_{18}$ 0.000 0.000 0.860 0.861 -1.426e-06 -1.423e-06 0.99 -1.634e-06 1.114e-06 0.99007
- $X_{19}$ 0.000 0.000 0.815 0.816 -1.517e-06 -1.516e-06 0.99 -1.716e-06 1.333e-06 0.99007
Table 3: Regression model feature importance ranking according to the empirical median (6) for one of the independent training/testing experiments that was randomly chosen. The table also reports the endpoints of the randomized CI (7) for the median at a $99\%$ level, the standard (non-randomized) two-sided CI (9) along with coverage probabilities, and the distribution of the randomized $p$-value.
The $p$-value distributions show that our approach is able to clearly distinguish between the significant and irrelevant features. The distributions of the significant features $\ell=1,\ldots,12$ are concentrated near 0, far below the level $\alpha=1\%$. This provides substantial evidence in favor of rejecting the null for these features (the rejection probability is equal to 1). The $p$-value distributions of the irrelevant features are concentrated significantly above $\alpha$, implying clear retention of the null for these features. There are no equivocal cases, eliminating the need to actually randomize.
Feature $\ell=9$ is judged to be the most important feature, followed by features $\ell=6,10$ and $12$. The remaining significant features are relatively less important compared with this group of variables. Features $\ell=4,5$ have (almost) identical medians and CIs and thus importance, consistent with their symmetric roles in the conditional mean function (5.1).
To demonstrate the stability of AICO feature importance scores, Figure 4 shows the empirical median (6) for each of the ten training/testing experiments. The AICO importance ranking is stable and robust to noise in the data. The ranking is also robust to alternative choices of the score function $S$ such as S(m,y)=|my|S(m,y)=-|m-y| (results available upon request). For comparison, Figure 4 also provides test-implied feature importance scores for CPT, HRT, and LOCO (?, ?, ?), kernel and deep SHAP feature importance (?), permutation feature importance scores (?, ?), and Gaussian and deep knockoff feature importance scores (?, ?). It is important to note that there is no ground truth for feature importance (?); hence we do not focus on the differences in importance rankings across alternative approaches. We do, however, highlight the robustness of the measures based on testing and variable selection procedures. The widely-used SHAP measures deliver inconsistent rankings across implementations (deep vs. kernel). The rankings are also highly unstable across experiments (data sets). That is, they appear to be highly susceptible to noise in the data. This questions the usefulness of SHAP for feature importance.
Figure 4: Feature importance scores across training/testing experiments.
Finally we consider the computational efficiency of our test vs. the aforementioned alternatives. Figure 5 reports average computation times across 10 experiments. [^11] The compute environment consists of 8GB memory and 4 CPU cores (Intel Xeon CPU E5-2670, 2.60GHz and Silver 4110 CPU, 2.10GHz). In each of these experiments we test (AICO, CPT, HRT, LOCO) or screen (knockoffs) all 19 features or, in the case of SHAP and permutation feature importance, compute importance scores for all 19 features. The time reported includes the time to train any sub or auxiliary models. It does not include the time to train the original neural network (14).
AICO achieves an unprecedented degree of computational efficiency. In terms of full-fledged testing procedures, AICO is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure. AICO is almost two orders of magnitudes more efficient than Gaussian knockoffs, the fastest screening procedure. It is more than three orders of magnitude faster than permutation importance, the fastest algorithmic feature importance measure. The testing results reported earlier in this section demonstrate that AICOs computational efficiency does not come at the expense of statistical performance.
Figure 5: Average computation times for testing/screening all 19 features.
5.2 Classification
We next consider a binary response $Y\in\{0,1\}$ governed by the model
$\mathbb{E}(Y|X)=\mathbb{P}(Y=1|X)=L(\mu(X))$ (16)
where $L(u)=1/(1+e^{u})$, the function $\mu$ is given by (5.1) and the features $X=(X^{1},\ldots,X^{d})$ are as defined in Section 5.1. We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. The data are imbalanced, as is often the case in practice (e.g. credit scoring): about $10\%$ of the samples have $Y_{i}=1$. We train a single-layer neural network model $f$ of the conditional mean $\mathbb{E}(Y|X)$:
$f(x)=L\left(b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x)\right)$ (17)
where $\psi(y)=1/(1+e^{-y})$, $K$ is the number of units, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters. The training procedure uses a standard cross-entropy criterion, $L_{2}$ regularization with weight $10^{-3}$ (approximately the optimal weight in the regression model training), Adam, a batch size of 32, and 500 epochs with early stopping. A validation set ($25\%$ of $I_{1}$) and a Bayesian approach are used to optimize the learning rate and the number of units, $K$. The optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$. The Area Under the ROC Curve on the training (and test) data exceeds $99\%$.
The testing procedure employs a score function $S(m,y)=y\log m+(1-y)\log(1-m)$ consistent with the training criterion. We use the baseline and comparison feature specification (5.1). Table 4 reports the results for confidence levels $\alpha=0.01,0.05$. As in the regression case, the test has perfect power and avoids false discoveries. The test is generally robust to alternative specifications of the baseline (e.g. using the median instead of the mean to construct $v^{\ell}$, setting $v^{\ell}$ to the training mean also for the 1-hot feature $\ell=9$). [^12] The local baseline has perfect power but some false discoveries. The test is also robust to using an alternative score function $S(m,y)=-(m-y)^{2}$.
$\ell$ AICO CPT HRT LOCO Deep Gaussian
Knockoff Knockoff
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 9 9 10 10 10 10 10 10
2 10 10 10 10 9 9 10 10 10 10 10 10
3 10 10 10 10 9 9 10 10 10 10 10 10
4 10 10 10 10 9 9 10 10 9 9 10 10
5 10 10 10 10 9 9 10 10 10 10 9 9
6 10 10 10 10 9 9 10 10 10 10 10 10
7 10 10 10 10 9 9 10 10 10 10 10 10
8 10 10 10 10 9 9 10 10 9 9 9 9
9 10 10 10 10 9 9 10 10 10 10 10 10
10 10 10 10 10 9 9 10 10 10 10 6 6
11 10 10 10 10 9 9 10 10 10 10 10 10
12 10 10 10 10 9 9 10 10 10 10 10 10
13 1 1 1 1 2 2 4 4 2 2 1 1
14 0 0 2 1 0 0 5 5 1 1 1 1
15 0 0 2 2 4 4 6 6 0 0 3 3
16 0 0 3 1 2 2 5 5 1 1 2 2
17 1 1 0 0 1 1 5 5 0 0 1 1
18 1 1 0 0 3 3 5 5 0 0 2 2
19 1 1 1 1 1 1 3 3 0 0 1 1
Table 4: Classification model test results for 10 independent training/testing experiments at two standard confidence levels ($\alpha=5\%$ and $\alpha=1\%$). The table compares AICO, LOCO, HRT, Deep Knockoff, and Gaussian Knockoff procedures, showing the total number of rejections across 10 trials. The test set size is $N=5\cdot 10^{5}$.
Table 4 also reports the results for LOCO, CPT, HRT and Gaussian and deep knockoff procedures. AICO, CPT and LOCO have full power. LOCO, however, suffers from significant number of false discoveries (which cannot be addressed with a standard Bonferroni adjustment). HRT and knockoff procedures have less power and some false discoveries. In this comparison, AICO outperforms the alternatives. It performs as well as in the regression setting.
AICO
$\ell$ $N=10^{5}$ $N=10^{4}$ $N=10^{3}$ $N=500$ $N=250$ $N=100$
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 10 10 10 10 9 7 3 0
2 10 10 10 10 6 5 3 2 6 5 3 1
3 10 10 10 10 10 10 10 10 10 10 8 8
4 10 10 10 10 10 10 10 9 6 2 4 0
5 10 10 10 10 10 10 8 8 8 6 6 1
6 10 10 10 10 10 10 10 10 8 6 2 1
7 10 10 8 7 3 3 2 0 1 0 0 0
8 10 10 5 2 0 0 0 0 0 0 0 0
9 10 10 10 10 6 3 3 1 5 3 3 1
10 10 10 10 10 10 10 10 10 9 9 8 5
11 10 10 10 10 6 2 9 5 0 0 3 1
12 10 10 10 10 10 9 6 5 5 2 2 0
13 0 0 1 0 1 0 3 0 0 0 0 0
14 0 0 0 0 0 0 1 0 1 0 2 1
15 0 0 0 0 0 0 0 0 1 1 1 0
16 0 0 0 0 0 0 0 0 0 0 2 0
17 1 1 1 0 1 0 2 0 0 0 0 0
18 1 1 1 0 1 0 2 0 0 0 0 0
19 0 0 1 0 1 0 0 0 0 0 0 0
Table 5: Classification model test results for 10 independent training/testing experiments, at two confidence levels ($\alpha=5\%$ and $\alpha=1\%$), across various test sample sizes $N$. We report the total number of rejections over 10 trials.
Table 5 reports test results for varying test set sample sizes $N$. Power decreases faster with $N$ than in the regression case, especially for features $\ell=7,8$. False discoveries appear to be fairly insensitive to $N$.
AICO, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 98.8% 1.2% coverage lower upper coverage
1 $X_{3}$ 3.577e-07 1.000 0.000 0.000 3.340e-07 3.340e-07 0.99 3.320e-07 3.829e-07 0.99007
2 $X_{10}$ 2.324e-08 1.000 0.000 0.000 2.077e-08 2.077e-08 0.99 2.053e-08 2.619e-08 0.99007
3 $X_{6}$ 6.722e-10 1.000 0.000 0.000 6.014e-10 6.016e-10 0.99 5.939e-10 7.656e-10 0.99007
4 $X_{1}$ 1.630e-10 1.000 0.000 0.000 1.442e-10 1.443e-10 0.99 1.426e-10 1.862e-10 0.99007
5 $X_{12}$ 6.989e-11 1.000 0.000 0.000 5.826e-11 5.828e-11 0.99 5.702e-11 8.597e-11 0.99007
6 $X_{5}$ 4.339e-11 1.000 0.000 0.000 3.697e-11 3.697e-11 0.99 3.643e-11 5.135e-11 0.99007
7 $X_{4}$ 3.169e-11 1.000 0.000 0.000 2.707e-11 2.707e-11 0.99 2.668e-11 3.733e-11 0.99007
8 $X_{11}$ 1.001e-11 1.000 0.000 0.000 8.027e-12 8.030e-12 0.99 7.821e-12 1.275e-11 0.99007
9 $X_{9}$ 6.787e-12 1.000 0.000 0.000 5.025e-12 5.025e-12 0.99 4.863e-12 9.439e-12 0.99007
10 $X_{2}$ 2.575e-14 1.000 0.000 0.000 1.823e-14 1.825e-14 0.99 1.755e-14 3.704e-14 0.99007
11 $X_{7}$ 1.100e-15 1.000 3.262e-71 3.431e-71 5.257e-16 5.261e-16 0.99 4.808e-16 2.326e-15 0.99007
12 $X_{8}$ 1.344e-17 1.000 1.706e-14 1.743e-14 2.386e-18 2.394e-18 0.99 1.972e-18 4.921e-17 0.99007
- $X_{13}$ 0.000 0.000 1.000 1.000 0.000 0.000 0.99 0.000 0.000 1.00000
- $X_{14}$ 0.000 0.000 1.000 1.000 -1.396e-21 -1.390e-21 0.99 -2.680e-21 0.000 0.99485
- $X_{15}$ -3.193e-21 0.000 1.000 1.000 -5.676e-20 -5.672e-20 0.99 -7.258e-20 0.000 0.99503
- $X_{16}$ 0.000 0.000 1.000 1.000 -3.066e-22 -3.052e-22 0.99 -5.372e-22 0.000 0.99501
- $X_{17}$ 0.000 0.000 1.000 1.000 0.000 0.000 0.99 0.000 0.000 0.99992
- $X_{18}$ 0.000 0.000 0.940 0.940 0.000 0.000 0.99 -1.849e-32 1.780e-22 0.99007
- $X_{19}$ 0.000 0.000 1.000 1.000 0.000 0.000 0.99 0.000 0.000 0.99999
Table 6: Classification model feature importance ranking according to the empirical median (6) for one of the independent training/testing experiments that was randomly chosen. The table also reports the endpoints of the randomized CI (7) for the median at a $99\%$ level, the standard (non-randomized) two-sided CI (9) along with coverage probabilities, and the distribution of the randomized $p$-value.
Table 6 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen. It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities. Table 3 also provides the distribution of the randomized $p$-value.
Figure 6 shows the empirical median (6) for each of the ten training/testing experiments, for each of several methods including AICO. Across methods, the scores of the less important variables tend to be (somewhat) more volatile than in the regression setting. SHAP delivers inconsistent and highly unstable rankings, as in the regression case.
Figure 6: Classification model feature importance across training/testing experiments.
Figure 7 reports average computation times across 10 experiments. As in the regression case, AICO achieves an unprecedented degree of computational efficiency. It is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure. AICO is an order of magnitude more efficient than Gaussian knockoffs, the fastest screening procedure. It is more than three orders of magnitude faster than permutation importance, the fastest algorithmic feature importance measure.
Figure 7: Average computation times for testing/screening all 19 features.
6 Empirical applications
6.1 Credit card default
This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.131313The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default. There are 23 explanatory variables (see Table 7). The dataset was split into training (60%), validation (15%), and testing (25%) sets with class proportions maintained through stratified sampling. Categorical variables are one-hot encoded and continuous variables are standardized (mean 0 and unit variance) based on training set statistics.
Feature Data Type Levels (for Categorical) Description
LIMIT_BAL Continuous Amount of the given credit (NT dollars).
SEX Categorical Male, Female Gender of the cardholder.
EDUCATION Categorical Graduate school, University, High school, Others Education level of the cardholder.
MARRIAGE Categorical Married, Single, Others Marital status of the cardholder.
AGE Continuous Age of the cardholder (in years).
[…]
PAY_2 Categorical Same as PAY_0 Repayment status in August 2005.
PAY_3 Categorical Same as PAY_0 Repayment status in July 2005.
PAY_4 Categorical Same as PAY_0 Repayment status in June 2005.
PAY_5 Categorical Same as PAY_0 Repayment status in May 2005.
PAY_6 Categorical Same as PAY_0 Repayment status in April 2005.
BILL_AMT1 Continuous Amount of bill statement (NT dollars) in September 2005.
BILL_AMT2 Continuous Amount of bill statement (NT dollars) in August 2005.
BILL_AMT3 Continuous Amount of bill statement (NT dollars) in July 2005.
BILL_AMT4 Continuous Amount of bill statement (NT dollars) in June 2005.
BILL_AMT5 Continuous Amount of bill statement (NT dollars) in May 2005.
BILL_AMT6 Continuous Amount of bill statement (NT dollars) in April 2005.
PAY_AMT1 Continuous Amount paid (NT dollars) in September 2005.
PAY_AMT2 Continuous Amount paid (NT dollars) in August 2005.
PAY_AMT3 Continuous Amount paid (NT dollars) in July 2005.
PAY_AMT4 Continuous Amount paid (NT dollars) in June 2005.
PAY_AMT5 Continuous Amount paid (NT dollars) in May 2005.
PAY_AMT6 Continuous Amount paid (NT dollars) in April 2005.
Table 7: Raw features of the credit default dataset.
We implement an XGBoost (?) classifier using a cross-entropy loss function. We use the area under the ROC curve (AUC) on the validation set as a metric for hyper-parameter optimization. We employ the Optuna package (?), utilizing the Tree-Structured Parzen Estimator (?), with 250 trials. The search space includes the boosting algorithm, tree-specific settings such as tree depth and minimum child weight, and regularization terms. Class imbalance was mitigated using a scaling weight parameter, and training was capped at 1,000 boosting rounds with early stopping if the validation AUC did not improve for 50 consecutive rounds. The optimal hyper-parameters include a gradient tree boosting algorithm (?) with an approximation-based tree-building strategy; an L2 regularization parameter (lambda) of 0.2 and an L1 regularization parameter (alpha) of 0.001; a subsampling ratio of 100% (no subsampling); a column sampling ratio for each tree of 65%; a maximum tree depth of six levels; a minimum child weight of two; a learning rate (eta) of 0.05; a minimum loss reduction (gamma) of 0.00004; and a loss-guided tree growth policy.
We train an ensemble of 100 XGBoost models on 100 bootstrap samples of the training data, all using the same hyper-parameters obtained during the tuning phase. Each model was trained for up to 1,000 boosting rounds, with early stopping triggered after 50 rounds of no improvement in the validation AUC. Ensemble predictions are obtained by averaging the outputs of all 100 models. The resulting ensemble model has AUCs of $0.837$ on the training set, $0.782$ on the validation set, and $0.786$ on the test set. (The results are robust with respect to the model. An ensemble of neural networks, for example, yields very similar results.)
As in the classification example of Section 5.2, we use a score function $S(m,y)=y\log m+(1-y)\log(1-m)$ to run the test. We continue to use the baseline and comparison feature specification (5.1), where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ in case feature $\ell$ is continuous. For a categorical variable we take the alternative mode $v_{i}^{\ell}=\arg\max_{x\in\mathcal{X}^{\ell}\setminus\{X_{i}^{\ell}\}}p^{\ell}(x|I_{1})$ where $\mathcal{X}^{\ell}$ is the (marginal) support of $X^{\ell}$ and $p^{\ell}(\cdot|I_{1})$ is the (empirical) probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$. This specification yields values in the support of a categorical. Here, we treat a categorical variable as a single feature rather than treating the associated one-hot variables separately. Alternatively, if higher resolution is required, we can run the test for each individual level of a categorical.
Table 8 summarizes the test results for $\alpha=1\%$ ($N=7,500$ throughout). The significant features, ranked according to the empirical median (6), are PAY_0 through PAY_6, MARRIAGE, SEX, and AGE. Past delinquencies are significant predictors of future payment default. The more recent the payment record the stronger is the influence of timely payment. The actual monthly bill and payment amounts (e.g., PAY_AMT and BILL_AMT) are not significant. Marital status, gender and age, but not education level, are significant predictors of default, even after controlling for past payment behavior. However, the influence of these variables is relatively small, especially for age (the associated medians are at least an order of magnitude smaller than the medians of the payment records). These results indicate that personal circumstances have a statistically significant influence on borrower behavior, beyond patterns in prior payment.
AICO, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 23.4% 76.6% coverage lower upper coverage
1 PAY_0 0.150 1.000 0.000 0.000 0.143 0.143 0.99 0.142 0.155 0.99063
2 PAY_2 0.121 1.000 0.000 0.000 0.117 0.117 0.99 0.116 0.125 0.99063
3 PAY_3 0.095 1.000 0.000 0.000 0.091 0.091 0.99 0.091 0.099 0.99063
4 PAY_4 0.059 1.000 0.000 0.000 0.055 0.055 0.99 0.054 0.064 0.99063
5 PAY_5 0.037 1.000 2.520e-322 6.423e-322 0.035 0.035 0.99 0.035 0.039 0.99063
6 PAY_6 0.030 1.000 1.020e-297 2.493e-297 0.029 0.029 0.99 0.029 0.032 0.99063
7 MARRIAGE 0.002 1.000 3.912e-11 4.561e-11 0.002 0.002 0.99 0.001 0.003 0.99063
8 SEX 0.002 1.000 1.026e-37 1.383e-37 0.001 0.002 0.99 0.001 0.002 0.99063
9 AGE 3.796e-04 1.000 1.788e-09 2.056e-09 2.200e-04 2.204e-04 0.99 2.036e-04 5.814e-04 0.99063
- LIMIT_BAL 0.000 0.000 0.794 0.800 0.000 0.000 0.99 0.000 3.603e-04 0.99496
- BILL_AMT1 -0.006 0.000 1.000 1.000 -0.007 -0.007 0.99 -0.008 -0.005 0.99063
- BILL_AMT2 -0.001 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -9.540e-04 0.99063
- BILL_AMT3 0.000 0.000 0.975 0.976 -1.599e-04 -1.592e-04 0.99 -1.751e-04 5.542e-05 0.99063
- BILL_AMT4 -4.094e-04 0.000 1.000 1.000 -5.298e-04 -5.284e-04 0.99 -5.504e-04 -2.773e-04 0.99063
- BILL_AMT5 -3.203e-04 0.000 1.000 1.000 -5.070e-04 -5.054e-04 0.99 -5.220e-04 -1.887e-04 0.99063
- BILL_AMT6 -1.176e-04 0.000 1.000 1.000 -2.041e-04 -2.023e-04 0.99 -2.160e-04 -3.890e-05 0.99063
- PAY_AMT1 -0.012 0.000 1.000 1.000 -0.014 -0.014 0.99 -0.014 -0.010 0.99063
- PAY_AMT2 -0.008 0.000 1.000 1.000 -0.009 -0.009 0.99 -0.010 -0.007 0.99063
- PAY_AMT3 -0.006 0.000 1.000 1.000 -0.006 -0.006 0.99 -0.006 -0.005 0.99063
- PAY_AMT4 -3.764e-04 0.000 1.000 1.000 -5.811e-04 -5.779e-04 0.99 -5.998e-04 -1.935e-04 0.99063
- PAY_AMT5 -0.002 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -0.002 0.99063
- PAY_AMT6 -0.001 0.000 1.000 1.000 -0.002 -0.001 0.99 -0.002 -0.001 0.99063
- EDUCATION -3.891e-04 0.000 0.848 0.853 -0.001 -0.001 0.99 -0.001 4.611e-04 0.99063
Table 8: AICO test summary for the credit card default dataset.
We can further explore the role of a feature by running the test on subsets of the test set. For example, Table 9 reports the results for a test on the 4,029 (out of 7,500) samples for which $\texttt{MARRIAGE}=\text{Single}$. While EDUCATION is not significant at the $1\%$ level, it is significant at the $5\%$ level (the distribution of the $p$-value is concentrated below $0.05$). Furthermore, EDUCATION is insignificant when running the test on the samples for which $\texttt{MARRIAGE}=\text{Married}$ (results available on request). This means educational level is a significant predictor of default for an unmarried borrower but not a married one.
AICO, Conditional on MARRIAGE = Single, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 83.0% 17.0% coverage lower upper coverage
1 PAY_0 0.151 1.000 1.833e-248 5.763e-248 0.143 0.143 0.99 0.142 0.158 0.99023
2 PAY_2 0.130 1.000 0.000 0.000 0.125 0.125 0.99 0.124 0.136 0.99023
3 PAY_3 0.100 1.000 1.072e-295 3.822e-295 0.095 0.095 0.99 0.095 0.105 0.99023
4 PAY_4 0.063 1.000 1.402e-260 4.555e-260 0.056 0.057 0.99 0.056 0.069 0.99023
5 PAY_5 0.040 1.000 4.099e-183 1.071e-182 0.037 0.037 0.99 0.037 0.044 0.99023
6 PAY_6 0.029 1.000 3.112e-171 7.848e-171 0.028 0.028 0.99 0.027 0.031 0.99023
7 AGE 0.002 1.000 2.601e-48 4.147e-48 0.002 0.002 0.99 0.002 0.002 0.99023
8 SEX 0.001 1.000 1.972e-16 2.560e-16 0.001 0.001 0.99 0.001 0.002 0.99023
- LIMIT_BAL -4.292e-04 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 0.000 0.99511
- BILL_AMT1 -0.005 0.000 1.000 1.000 -0.007 -0.007 0.99 -0.007 -0.003 0.99023
- BILL_AMT2 -0.001 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -8.918e-04 0.99023
- BILL_AMT3 4.366e-05 0.000 0.376 0.388 -1.597e-05 -1.038e-05 0.99 -3.777e-05 2.814e-04 0.99023
- BILL_AMT4 -4.339e-04 0.000 1.000 1.000 -6.109e-04 -6.105e-04 0.99 -6.237e-04 -2.882e-04 0.99023
- BILL_AMT5 -1.737e-04 0.000 1.000 1.000 -3.338e-04 -3.327e-04 0.99 -3.612e-04 -3.330e-05 0.99023
- BILL_AMT6 -1.147e-04 0.000 1.000 1.000 -2.326e-04 -2.315e-04 0.99 -2.509e-04 0.000 0.99474
- PAY_AMT1 -0.014 0.000 1.000 1.000 -0.015 -0.015 0.99 -0.016 -0.011 0.99023
- PAY_AMT2 -0.010 0.000 1.000 1.000 -0.012 -0.012 0.99 -0.012 -0.009 0.99023
- PAY_AMT3 -0.006 0.000 1.000 1.000 -0.008 -0.008 0.99 -0.008 -0.005 0.99023
- PAY_AMT4 -4.392e-04 0.000 1.000 1.000 -7.181e-04 -7.118e-04 0.99 -7.488e-04 -1.825e-04 0.99023
- PAY_AMT5 -0.002 0.000 1.000 1.000 -0.003 -0.003 0.99 -0.003 -0.002 0.99023
- PAY_AMT6 -0.002 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -0.001 0.99023
- EDUCATION 0.001 0.000 0.012 0.013 -7.817e-05 -4.987e-05 0.99 -2.034e-04 0.002 0.99023
Table 9: AICO test summary for the credit card default dataset. The test is conditioned the samples for which $\texttt{MARRIAGE}=\text{Single}$.
In the US, marital status, gender and age are protected Characteristics under Fair Lending laws, so these variables are not used (they might not even be collected) to construct default probability estimates or credit scores for purposes of credit underwriting. Our results indicate, within the context of Taiwanese borrowers, that these attributes represent signals useful for risk prediction. All else equal, a score constructed without this set of variables might be less accurate.
6.2 Mortgage risk
We next consider AICO feature significance for a more complex classification model trained to predict the state of a mortgage loan in the next month given a history of a time-varying feature vector available at the beginning of the month. The 7 states are Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent, Foreclosure, Paid Off, and Real Estate Owned. The analysis is based on an extensive panel dataset of mortgages from 4 ZIP codes in the greater Los Angeles area which are tracked monthly between 1/1994 and 12/2022 (this is a subset of the commercial CoreLogic Loan-Level Market Analytics dataset). The feature vector is of dimension $d=52$, and includes variables such as the current state, FICO credit score, loan balance, current interest rate, market interest rate, unemployment rate, etc., many of which are time-varying.
We use the data filtering procedure of ?, resulting in roughly 5 million loan-month transitions from 117,523 active loans. The train/validation/test splits are as follows: 1/1994 - 6/2009, 7/2009 - 12/2009, and 1/2010 - 12/2022. We adopt the set-sequence model of ? with LongConv (?) for the sequence layer, which represents the current state of the art for mortgage transition data. We closely follow the training, validation and testing approach of ?.
We use the feature effect grouping approach discussed in Section 2.2 for panel data, where a unit is a mortgage loan. The panel is unbalanced because loans enter the dataset at different times and are observed for different time periods. The negative multi-class cross-entropy loss is used as the score function $S$, leading to a test of feature importance jointly for transitions between all states. Alternative formulations are possible; for example, we could target specific transitions (e.g. current to prepayment). Baseline and comparison features are chosen as in Section 6.1, at the level of a group.
Table 10 summarizes the test results for $\alpha=1\%$ ($N=46,800$). In line with expectations, the most important feature by far is the current state of the mortgage. The market mortgage rate is the second most important feature, highlighting the significance of time-varying macro-economic conditions for mortgage borrower behavior. Indeed, borrowers have an incentive to prepay and refinance their mortgage if the market mortgage rate falls below their own current loan interest rate (the latter is ranked eighth). We also recognize path-dependency: the number of times the borrower was current and the number of times the borrower was 30 days delinquent during the past 12 months are highly significant and economically important variables for borrower behavior.
AICO, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 37.5% 62.5% coverage lower upper coverage
1 current_state 0.560 1.000 0.000 0.000 0.557 0.557 0.99 0.556 0.564 0.99016
2 national_mortgage_rate 0.006 1.000 5.372e-66 6.297e-66 0.005 0.005 0.99 0.005 0.006 0.99016
3 pool_insurance_flag_value 0.005 1.000 4.345e-314 6.182e-314 0.005 0.005 0.99 0.005 0.005 0.99016
4 times_current 0.002 1.000 5.920e-178 7.708e-178 0.002 0.002 0.99 0.002 0.003 0.99016
5 times_30dd_value 0.002 1.000 0.000 0.000 0.001 0.001 0.99 0.001 0.002 0.99029
6 loan_age 7.823e-04 1.000 3.830e-18 4.151e-18 5.401e-04 5.407e-04 0.99 5.230e-04 0.001 0.99016
7 lagged_prepayment_rate 7.518e-04 1.000 7.103e-127 8.870e-127 6.730e-04 6.732e-04 0.99 6.662e-04 8.298e-04 0.99029
8 current_interest_rate_value 6.497e-04 1.000 3.066e-83 3.667e-83 5.727e-04 5.730e-04 0.99 5.636e-04 7.351e-04 0.99016
9 inferred_collateral_type_value 5.229e-04 1.000 2.792e-43 3.173e-43 4.318e-04 4.318e-04 0.99 4.231e-04 6.206e-04 0.99016
10 mba_days_delinquent_value 3.738e-04 1.000 1.890e-131 2.370e-131 3.379e-04 3.380e-04 0.99 3.341e-04 4.136e-04 0.99016
11 initial_interest_rate 2.199e-04 1.000 1.219e-27 1.349e-27 1.775e-04 1.776e-04 0.99 1.751e-04 2.671e-04 0.99016
12 times_foreclosure 2.032e-04 1.000 2.492e-46 2.844e-46 1.680e-04 1.681e-04 0.99 1.631e-04 2.454e-04 0.99029
13 original_balance 1.815e-04 1.000 4.651e-43 5.283e-43 1.505e-04 1.505e-04 0.99 1.479e-04 2.178e-04 0.99016
14 prepay_penalty_flag_value 1.568e-04 1.000 6.926e-19 7.520e-19 1.152e-04 1.152e-04 0.99 1.115e-04 2.001e-04 0.99016
15 buydown_flag_value 1.294e-04 1.000 3.980e-35 4.462e-35 1.062e-04 1.064e-04 0.99 1.035e-04 1.536e-04 0.99016
16 fico_score_at_origination 1.233e-04 1.000 4.995e-36 5.609e-36 9.650e-05 9.669e-05 0.99 9.399e-05 1.515e-04 0.99016
17 scheduled_monthly_pi_value 3.479e-05 1.000 2.451e-30 2.726e-30 2.661e-05 2.663e-05 0.99 2.576e-05 4.341e-05 0.99016
- original_ltv 1.874e-05 0.000 0.017 0.018 -2.449e-06 -2.418e-06 0.99 -4.904e-06 4.245e-05 0.99016
- unemployment_rate -3.907e-04 0.000 1.000 1.000 -4.822e-04 -4.813e-04 0.99 -4.909e-04 -2.838e-04 0.99016
- current_balance_value -1.298e-06 0.000 0.520 0.524 -3.206e-05 -3.195e-05 0.99 -3.507e-05 3.639e-05 0.99016
- scheduled_principal_value -7.242e-06 0.000 1.000 1.000 -1.134e-05 -1.133e-05 0.99 -1.176e-05 -3.263e-06 0.99016
- convertible_flag_value -1.886e-04 0.000 1.000 1.000 -2.477e-04 -2.476e-04 0.99 -2.533e-04 -1.270e-04 0.99016
- io_flag_value -2.458e-04 0.000 1.000 1.000 -2.948e-04 -2.946e-04 0.99 -2.989e-04 -2.044e-04 0.99016
- negative_amortization_flag_value 9.342e-05 0.000 0.015 0.015 -7.297e-06 -6.803e-06 0.99 -1.626e-05 2.213e-04 0.99016
- original_term_value -1.047e-04 0.000 0.999 0.999 -1.854e-04 -1.852e-04 0.99 -1.936e-04 -1.765e-05 0.99016
- times_60dd -3.928e-04 0.000 1.000 1.000 -5.032e-04 -5.024e-04 0.99 -5.118e-04 -2.708e-04 0.99016
- times_90dd -4.781e-04 0.000 1.000 1.000 -5.630e-04 -5.630e-04 0.99 -5.723e-04 -3.823e-04 0.99016
- lagged_foreclosure_rate -1.049e-04 0.000 0.998 0.998 -1.769e-04 -1.765e-04 0.99 -1.846e-04 -9.468e-06 0.99016
Table 10: AICO test summary for the mortgage risk dataset with set-sequence model. Training data is now grouped and treated at sample-loan level. Continuous features are aggregated with mean while discrete and categorical features are aggregated by mode (just basic mode).
7 Conclusion
This paper develops a simple, statistically rigorous, and computationally efficient framework for assessing feature significance in supervised learning. Our randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation. Empirical studies show strong performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, demonstrating its ability to inform high-stakes decisions with legal, economic, and regulatory implications. More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines rigor with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
Appendix A Proofs of technical results
Proof of Proposition 3.2.
Under Assumption 3.1, $n_{+}(x)\sim B_{N,\theta_{x}}$ where $\theta_{x}=1-F_{\Delta}(x)$. The Binomial distribution $B_{N,\theta_{x}}$ has a monotone likelihood ratio, so the Neyman-Pearson lemma (?, Theorem 3.4.1) guarantees the existence of a UMP test for the hypotheses
$\theta_{x}\leq\theta_{0}\quad\text{vs.}\quad\theta_{x}>\theta_{0}$
for some $\theta_{0}\in[0,1]$. Due to the continuity of $F_{\Delta}$, we have $F_{\Delta}(M)=1/2$ so for $\theta_{0}=1/2$ and $x=M_{0}$ these hypotheses are equivalent to (1). The UMP test has test function $\phi(n)=1\{n>T\}+\gamma 1\{n=T\}$ for some $T\in\mathbb{N}$ and $\gamma\in[0,1]$ determined by the requirement that the test’s size be equal to $\alpha$. Formally, for any $M_{0}\in\mathbb{R}$,
$\alpha$ $=\mathbb{E}_{M_{0}}(\phi(n_{+}(M_{0})))$
$=\mathbb{P}_{M_{0}}(n_{+}(M_{0})>T)+\mathbb{E}_{M_{0}}(\gamma|n_{+}(M_{0})=T)\mathbb{P}_{M_{0}}(n_{+}(M_{0})=T).$
It follows that $T=T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ and
$\gamma=\gamma_{N}(T,\alpha)=\frac{\alpha-\mathbb{P}_{M_{0}}(n_{+}(M_{0})>T)}{\mathbb{P}_{M_{0}}(n_{+}(M_{0})=T)}.$
We clearly have $n_{+}(M_{0})\sim B_{N,1/2}$ under $\mathbb{P}_{M_{0}}$ so $\mathbb{P}_{M_{0}}(n_{+}(M_{0})\leq n)=\Pi_{N}(n)$ and furthermore, $\mathbb{P}_{M_{0}}(n_{+}(M_{0})=n)=\pi_{N}(n)$. This gives (3). ∎
Proof of Proposition 4.1.
Let $U\sim U(0,1)$ and $I(x)=1-\phi_{N}(n_{+}(x),\alpha)$, where $n_{+}(x)=\#\{i\in I_{2}:\Delta_{i}>x\}$ for $x\in\mathbb{R}$, as in Proposition 3.2. From ?, a randomized CI for $M$ is given by the $U$-cut of $I(x)$:
$C_{N}(\alpha;U)=\{x\in\mathbb{R}\;|\;I(x)\geq U\}.$
Then, by iterated expectations and the independence of $U$, we get
$\mathbb{P}_{M}(M\in C_{N}(\alpha;U))$ $=\mathbb{P}_{M}(U\leq I(M))$
$=\mathbb{E}_{M}(1-\phi_{N}(n_{+}(M),\alpha))$
$=1-\alpha$
by (4). (Recall that $n_{+}(M)\sim B_{N,1/2}$ under $\mathbb{P}_{M}$.)
Since $n_{+}(M)=N-i$ for $M\in[\Delta_{(i)},\Delta_{(i+1)})$, we have
$C_{N}(\alpha;U)=\{[\Delta_{(i)},\Delta_{(i+1)})\;|\;i=0,\ldots,N;\;1-\phi_{N}(N-i,\alpha)\geq U\}.$
Moreover, note that
$1-\phi_{N}(N-i,\alpha)$ $=1\{N-i\leq T_{N,\alpha}\}-\gamma_{N}(T_{N,\alpha},\alpha)1\{N-i=T_{N,\alpha}\}$
$=1\{N-T_{N,\alpha}<i\}+(1-\gamma_{N}(T_{N,\alpha},\alpha))1\{N-T_{N,\alpha}=i\}.$
Thus $C_{N}(\alpha;U)$ is equal to the union of $\Delta_{(N-T_{N,\alpha})},\Delta_{(N-T_{N,\alpha}+1)})$ and $\Delta_{(N-T_{N,\alpha}+1)},\infty)$ if $1-\gamma_{N}(T_{N,\alpha},\alpha)\geq U$ and $\Delta_{(N-T_{N,\alpha}+1)},\infty)$ otherwise. ∎
Proof of Proposition 4.2.
Conditional independence implies that
$\mu(X)=\mathbb{E}(Y|X)=\mathbb{E}(Y|X^{-1})=g(X^{-1})$
for some function $g$, so $\mu(\cdot)$ does not depend on the first entry of its argument vector. By construction, $\overline{X}$ only differs from $\underline{X}$ at the first entry. Therefore $\mu(\underline{X})=\mu(\overline{X})$ for any pair $(\underline{X},\overline{X})$, and hence the feature effect is identically zero. ∎
Proof of Proposition 4.3.
Letting $\beta^{-1}=(\beta^{2},\ldots,\beta^{d})$, under the stated hypotheses we get
$\mathbb{E}(Y|X^{1})$ $=\beta\cdot\mathbb{E}(X|X^{1})+\mathbb{E}(\epsilon|X^{1})$
$=\beta^{1}X^{1}+\beta^{-1}\cdot\mathbb{E}(X^{-1})$
$=\mu(\overline{X}).$
With the $L_{2}$ score function,
$\Delta$ $=(Y-\mu(\underline{X}))^{2}-(Y-\mu(\overline{X}))^{2}$
$=\left((Y-\mu(\underline{X}))-(Y-\mu(\overline{X}))\right)\left((Y-\mu(\underline{X}))+(Y-\mu(\overline{X}))\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(\mu(\underline{X})+\mu(\overline{X})-2Y\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(\mu(\underline{X})-\mu(\overline{X})-2(Y-\mu(\overline{X}))\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}-2\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(Y-\mu(\overline{X})\right).$
Since $\underline{X}$ is deterministic, and $X^{1}$ is the only source of randomness in $\overline{X}$,
$\mathbb{E}(\Delta|X^{1})$ $=\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}-2\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(\mathbb{E}(Y|X^{1})-\mu(\overline{X})\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}.$
By iterated expectations, we conclude that
$\mathbb{E}(\Delta)$ $=\mathbb{E}\left(\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}\right)$
$=(\beta^{1})^{2}\mathbb{E}\left(\left(\mathbb{E}(X^{1})-X^{1}\right)^{2}\right),$
which completes the proof. ∎
Proof of Proposition $4.$
From Proposition 4.3, we must have that $\mathbb{E}(\Delta)=(\beta^{1})^{2}var(X^{1})=0$, which implies that $\beta^{1}=0$ when $X^{1}$ has strictly positive variance. Therefore,
$Y=\beta^{-1}\cdot X^{-1}+\epsilon.$
The independence of $\epsilon$ and $X$ then implies conditional independence. ∎

### v3

@@
 Alternatively, we might select an appropriate baseline value from the set $\{X_{k}:k\in I_{1}\}$, namely an element that is near the raw off-manifold value emerging from the baseline selection procedure.
-For clarity, the discussion above has focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$.
+[^6] We also allow for the selection and aggregation of alternative baselines for a given sample.
+A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
+2.2 Multiple features, panel data and missing data
+We have focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$.
 The approach naturally extends to testing multiple features, i.e.
 subsets of $[d]$, which facilitates the exploration of higher-order effects including feature interactions.
-We also allow for the selection and aggregation of alternative baselines for a given sample.
-A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
-2.2 Panel data
 Panel data settings, where a collection of units are observed over some time period, are prevalent in many domains.
-While our approach assumes independent samples, a simple extension—grouping—allow us to treat some dependent settings like panel data.
-The idea is to aggregate feature effects within each unit of dependence, so that dependence is contained within groups with inter-group dependencies being mitigated.
-Concretely, consider panel data $\{Z_{i,t}\}$, where $i=1,\dots,m$ indexes units and $t=1,\dots,T$ indexes time periods.
-We take as given a unit-level random split of the index set $m]=\{1,\dots,m\}$ into disjoint training and testing subsets $I_{1}$ and $I_{2}$, respectively, such that each unit, along with its entire time series, is assigned exclusively to either $I_{1}$ or $I_{2}$.
-We treat the test set $I_{2}$ as a collection of $|I_{2}|T$ samples and compute the per-unit-time feature effects $\Delta_{i,t}$ for each $(i,t)\in I_{2}$.
-We then aggregate these effects within each unit: $\Delta_{i}=\frac{1}{T}\sum_{t=1}^{T}\Delta_{i,t}$.
-For purposes of computing baseline and comparison features in this group setting, we also treat the training data $I_{1}$ at a unit level.
-For each $(i,t)\in I_{1}$ and feature $j\in[d]$, the per-unit-time samples are aggregated to obtain grouped feature samples $X_{i}^{j}=a_{i}^{j}(X^{j}_{i,1},\dots,X^{j}_{i,T})$, where $a_{i}^{j}$ is an aggregation function for the time series associated with unit $i$ (e.g.
-the mean, median, mode).
-The grouping approach naturally extends to unbalanced panel data, where subjects may be observed at different time points or for different durations.
-In such cases, the average for each subject is computed only over the observed periods.
-More broadly, the grouping approach can be adapted to other dependence structures beyond temporal ones.
-For instance, for spatially dependent data, one could group feature effects by region or neighborhood, under the assumption that neighborhoods are (approximately) independent.
-Similarly, in settings with repeated measurements, grouping by the underlying entity or context is effective.
+Appendix B discusses two simple extensions—conditioning and grouping—to construct the feature effect for panel data.
+Real-world data sets often have missing feature values. Appendix C shows how to treat missing or imputed values when constructing $\Delta_{i}$.
 3 Hypothesis testing
@@
 We run 10 independent training/testing experiments to assess the behavior of the test.
-The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.888Limited numerical precision can lead to ties, i.e.
+The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.999Limited numerical precision can lead to ties, i.e.
 samples $\Delta_{i}$ whose true values are distinct are being rounded to the same value.
@@
 [^11] Ties of samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing CIs and coverage probabilities.
-This is consistent with their treatment in the testing procedure (footnote 8).
+This is consistent with their treatment in the testing procedure (footnote 9).
 AICO, Level $\alpha=1\%$
@@
 6.1 Credit card default
-This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.131313The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default.
+This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.141414The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default.
 There are 23 explanatory variables (see Table 7).
@@
 We closely follow the training, validation and testing approach of ?.
-We use the feature effect grouping approach discussed in Section 2.2 for panel data, where a unit is a mortgage loan.
-The panel is unbalanced because loans enter the dataset at different times and are observed for different time periods.
-The negative multi-class cross-entropy loss is used as the score function $S$, leading to a test of feature importance jointly for transitions between all states.
-Alternative formulations are possible; for example, we could target specific transitions (e.g.
-current to prepayment).
-Baseline and comparison features are chosen as in Section 6.1, at the level of a group.
-Table 10 summarizes the test results for $\alpha=1\%$ ($N=46,800$).
-In line with expectations, the most important feature by far is the current state of the mortgage.
-The market mortgage rate is the second most important feature, highlighting the significance of time-varying macro-economic conditions for mortgage borrower behavior.
-Indeed, borrowers have an incentive to prepay and refinance their mortgage if the market mortgage rate falls below their own current loan interest rate (the latter is ranked eighth).
-We also recognize path-dependency: the number of times the borrower was current and the number of times the borrower was 30 days delinquent during the past 12 months are highly significant and economically important variables for borrower behavior.
+The dataset contains missing feature values, which are handled using missingness dummies and forward-filling based on the last observed value for each loan.
+To avoid biased test results, loan–month samples with missing values are excluded from the test set, as discussed in Appendix C.
+Because the extent of missingness varies across features, each feature test is conducted on a different effective sample size.
+We apply the feature-effect grouping approach described in Appendix B for panel data, treating each mortgage loan as a unit.
+The panel is unbalanced because loans enter the dataset at different times and are observed for varying durations.
+The score function $S$ is the negative mean multi-class cross-entropy loss, where the mean is computed across time steps for each loan.
+This multi-period specification of $S$ enables a joint test of feature importance across all state transitions.
+Alternative formulations are also possible—for instance, targeting specific transitions such as from Current to Prepaid.
+Baseline features are constructed through a two-step aggregation process: first from the loan–month level to the loan level, and then across all loans.
+For each feature $\ell$ and each loan $k\in I_{1}$ in the training set, the loan-level statistic $u_{k}^{\ell}$ is computed from the loan’s training records $\{X_{k,\cdot}:(k,\cdot)\in I_{1}\}$.
+Continuous variables are aggregated using the mean, whereas discrete and categorical variables use the mode.
+Using these loan-level aggregates, baseline and comparison features are then constructed according to (5.1), where $v_{i}^{\ell}$ is the mean of $\{u_{k}^{\ell}:k\in I_{1}\}$ for continuous features.
+For discrete and categorical variables, we instead use the alternative mode $v_{i}^{\ell}=\arg\max_{u\in\mathcal{X}^{\ell}\setminus\{u_{i}^{\ell}\}}p^{\ell}(u\mid I_{1})$, where $\mathcal{X}^{\ell}$ denotes the marginal support of $X^{\ell}$ and $p^{\ell}(\cdot\mid I_{1})$ is the empirical probability mass function of $\{u_{k}^{\ell}:k\in I_{1}\}$.
+This two-step aggregation ensures uniform weighting across loans when constructing baseline values, regardless of the number of observations available for each loan.
+Table 10 summarizes the test results for $\alpha=1\%$.
+As expected, the most important feature by far is the current state of the mortgage.
+The market mortgage rate is the second most important feature, underscoring the influence of time-varying macroeconomic conditions on borrower behavior.
+Borrowers have a strong incentive to prepay or refinance their mortgage when the market rate falls below their own loan rate, which is ranked nineth in importance.
+The results also reveal pronounced path dependence: the number of months a borrower remained current and the number of months 30 days delinquent during the past 12 months are both highly significant and economically important predictors of borrower behavior.
 AICO, Level $\alpha=1\%$
 P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
-rank variable median P(reject) lower upper 37.5% 62.5% coverage lower upper coverage
-1 current_state 0.560 1.000 0.000 0.000 0.557 0.557 0.99 0.556 0.564 0.99016
-2 national_mortgage_rate 0.006 1.000 5.372e-66 6.297e-66 0.005 0.005 0.99 0.005 0.006 0.99016
-3 pool_insurance_flag_value 0.005 1.000 4.345e-314 6.182e-314 0.005 0.005 0.99 0.005 0.005 0.99016
-4 times_current 0.002 1.000 5.920e-178 7.708e-178 0.002 0.002 0.99 0.002 0.003 0.99016
-5 times_30dd_value 0.002 1.000 0.000 0.000 0.001 0.001 0.99 0.001 0.002 0.99029
-6 loan_age 7.823e-04 1.000 3.830e-18 4.151e-18 5.401e-04 5.407e-04 0.99 5.230e-04 0.001 0.99016
-7 lagged_prepayment_rate 7.518e-04 1.000 7.103e-127 8.870e-127 6.730e-04 6.732e-04 0.99 6.662e-04 8.298e-04 0.99029
-8 current_interest_rate_value 6.497e-04 1.000 3.066e-83 3.667e-83 5.727e-04 5.730e-04 0.99 5.636e-04 7.351e-04 0.99016
-9 inferred_collateral_type_value 5.229e-04 1.000 2.792e-43 3.173e-43 4.318e-04 4.318e-04 0.99 4.231e-04 6.206e-04 0.99016
-10 mba_days_delinquent_value 3.738e-04 1.000 1.890e-131 2.370e-131 3.379e-04 3.380e-04 0.99 3.341e-04 4.136e-04 0.99016
-11 initial_interest_rate 2.199e-04 1.000 1.219e-27 1.349e-27 1.775e-04 1.776e-04 0.99 1.751e-04 2.671e-04 0.99016
-12 times_foreclosure 2.032e-04 1.000 2.492e-46 2.844e-46 1.680e-04 1.681e-04 0.99 1.631e-04 2.454e-04 0.99029
-13 original_balance 1.815e-04 1.000 4.651e-43 5.283e-43 1.505e-04 1.505e-04 0.99 1.479e-04 2.178e-04 0.99016
-14 prepay_penalty_flag_value 1.568e-04 1.000 6.926e-19 7.520e-19 1.152e-04 1.152e-04 0.99 1.115e-04 2.001e-04 0.99016
-15 buydown_flag_value 1.294e-04 1.000 3.980e-35 4.462e-35 1.062e-04 1.064e-04 0.99 1.035e-04 1.536e-04 0.99016
-16 fico_score_at_origination 1.233e-04 1.000 4.995e-36 5.609e-36 9.650e-05 9.669e-05 0.99 9.399e-05 1.515e-04 0.99016
-17 scheduled_monthly_pi_value 3.479e-05 1.000 2.451e-30 2.726e-30 2.661e-05 2.663e-05 0.99 2.576e-05 4.341e-05 0.99016
-- original_ltv 1.874e-05 0.000 0.017 0.018 -2.449e-06 -2.418e-06 0.99 -4.904e-06 4.245e-05 0.99016
-- unemployment_rate -3.907e-04 0.000 1.000 1.000 -4.822e-04 -4.813e-04 0.99 -4.909e-04 -2.838e-04 0.99016
-- current_balance_value -1.298e-06 0.000 0.520 0.524 -3.206e-05 -3.195e-05 0.99 -3.507e-05 3.639e-05 0.99016
-- scheduled_principal_value -7.242e-06 0.000 1.000 1.000 -1.134e-05 -1.133e-05 0.99 -1.176e-05 -3.263e-06 0.99016
-- convertible_flag_value -1.886e-04 0.000 1.000 1.000 -2.477e-04 -2.476e-04 0.99 -2.533e-04 -1.270e-04 0.99016
-- io_flag_value -2.458e-04 0.000 1.000 1.000 -2.948e-04 -2.946e-04 0.99 -2.989e-04 -2.044e-04 0.99016
-- negative_amortization_flag_value 9.342e-05 0.000 0.015 0.015 -7.297e-06 -6.803e-06 0.99 -1.626e-05 2.213e-04 0.99016
-- original_term_value -1.047e-04 0.000 0.999 0.999 -1.854e-04 -1.852e-04 0.99 -1.936e-04 -1.765e-05 0.99016
-- times_60dd -3.928e-04 0.000 1.000 1.000 -5.032e-04 -5.024e-04 0.99 -5.118e-04 -2.708e-04 0.99016
-- times_90dd -4.781e-04 0.000 1.000 1.000 -5.630e-04 -5.630e-04 0.99 -5.723e-04 -3.823e-04 0.99016
-- lagged_foreclosure_rate -1.049e-04 0.000 0.998 0.998 -1.769e-04 -1.765e-04 0.99 -1.846e-04 -9.468e-06 0.99016
-Table 10: AICO test summary for the mortgage risk dataset with set-sequence model.
-Training data is now grouped and treated at sample-loan level.
-Continuous features are aggregated with mean while discrete and categorical features are aggregated by mode (just basic mode).
+rank variable median P(reject) lower upper lower-1 lower-2 P(lower-2) coverage lower upper coverage missing rate
+1 current_state 0.560 1.000 0.000 0.000 0.557 0.557 0.625 0.99 0.556 0.564 0.9902 0.000
+2 national_mortgage_rate 0.006 1.000 5.372e-66 6.297e-66 0.005 0.005 0.625 0.99 0.005 0.006 0.9902 0.000
+3 pool_insurance_flag_value 0.005 1.000 1.031e-240 1.560e-240 0.005 0.005 0.032 0.99 0.005 0.006 0.9903 0.443
+4 times_current 0.002 1.000 5.920e-178 7.708e-178 0.002 0.002 0.625 0.99 0.002 0.003 0.9902 0.000
+5 times_30dd_value 0.002 1.000 0.000 0.000 0.001 0.001 0.393 0.99 0.001 0.002 0.9902 0.014
+6 loan_age 7.823e-04 1.000 3.830e-18 4.151e-18 5.401e-04 5.407e-04 0.625 0.99 5.230e-04 0.001 0.9902 0.000
+7 inferred_collateral_type_value 7.683e-04 1.000 2.425e-70 2.882e-70 6.749e-04 6.750e-04 0.535 0.99 6.652e-04 8.658e-04 0.9902 0.096
+8 lagged_prepayment_rate 7.518e-04 1.000 7.103e-127 8.870e-127 6.730e-04 6.732e-04 0.625 0.99 6.662e-04 8.298e-04 0.9903 0.000
+9 current_interest_rate_value 6.571e-04 1.000 8.549e-85 1.024e-84 5.789e-04 5.792e-04 0.530 0.99 5.716e-04 7.453e-04 0.9902 0.003
+10 mba_days_delinquent_value 3.738e-04 1.000 1.890e-131 2.370e-131 3.379e-04 3.380e-04 0.625 0.99 3.341e-04 4.136e-04 0.9902 0.000
+11 prepay_penalty_flag_value 2.733e-04 1.000 4.209e-56 4.903e-56 2.365e-04 2.366e-04 0.078 0.99 2.317e-04 3.184e-04 0.9901 0.085
+12 initial_interest_rate 2.199e-04 1.000 1.219e-27 1.349e-27 1.775e-04 1.776e-04 0.625 0.99 1.751e-04 2.671e-04 0.9902 0.000
+13 buydown_flag_value 2.192e-04 1.000 1.214e-68 1.461e-68 1.882e-04 1.883e-04 0.160 0.99 1.856e-04 2.537e-04 0.9903 0.235
+14 times_foreclosure 2.032e-04 1.000 2.492e-46 2.844e-46 1.680e-04 1.681e-04 0.625 0.99 1.631e-04 2.454e-04 0.9903 0.000
+15 original_balance 1.815e-04 1.000 4.651e-43 5.283e-43 1.505e-04 1.505e-04 0.625 0.99 1.479e-04 2.178e-04 0.9902 0.000
+16 fico_score_at_origination 1.233e-04 1.000 4.995e-36 5.609e-36 9.650e-05 9.669e-05 0.625 0.99 9.399e-05 1.515e-04 0.9902 0.000
+17 scheduled_monthly_pi_value 3.616e-05 1.000 4.464e-32 4.979e-32 2.805e-05 2.809e-05 0.067 0.99 2.688e-05 4.472e-05 0.9900 0.003
+- original_ltv 1.874e-05 0.000 0.017 0.018 -2.449e-06 -2.418e-06 0.625 0.99 -4.904e-06 4.245e-05 0.9902 0.000
+- unemployment_rate -3.907e-04 0.000 1.000 1.000 -4.822e-04 -4.813e-04 0.625 0.99 -4.909e-04 -2.838e-04 0.9902 0.000
+- current_balance_value -1.298e-06 0.000 0.520 0.524 -3.206e-05 -3.195e-05 0.625 0.99 -3.507e-05 3.639e-05 0.9902 0.000
+- scheduled_principal_value -5.314e-05 0.000 1.000 1.000 -5.989e-05 -5.983e-05 0.980 0.99 -6.061e-05 -4.456e-05 0.9901 0.573
+- convertible_flag_value -5.164e-04 0.000 1.000 1.000 -6.553e-04 -6.551e-04 0.034 0.99 -6.722e-04 -3.525e-04 0.9900 0.555
+- io_flag_value -4.553e-05 0.000 0.973 0.973 -1.026e-04 -1.020e-04 0.868 0.99 -1.091e-04 1.431e-05 0.9902 0.324
+- negative_amortization_flag_value -8.336e-04 0.000 1.000 1.000 -0.001 -0.001 0.825 0.99 -0.001 -4.979e-04 0.9903 0.853
+- original_term_value -1.041e-04 0.000 0.999 0.999 -1.843e-04 -1.843e-04 0.139 0.99 -1.924e-04 -1.752e-05 0.9900 1.066e-04
+- times_60dd -3.928e-04 0.000 1.000 1.000 -5.032e-04 -5.024e-04 0.625 0.99 -5.118e-04 -2.708e-04 0.9902 0.000
+- times_90dd -4.781e-04 0.000 1.000 1.000 -5.630e-04 -5.630e-04 0.625 0.99 -5.723e-04 -3.823e-04 0.9902 0.000
+- lagged_foreclosure_rate -1.049e-04 0.000 0.998 0.998 -1.769e-04 -1.765e-04 0.625 0.99 -1.846e-04 -9.468e-06 0.9902 0.000
+Table 10: AICO test summary for the mortgage risk dataset using the set-sequence model.
+Training data is grouped and treated at the loan-month level.
+The missing rate denotes the ratio of feature effects dropped from the test set due to missigness to the total number of loan-level feature effects after grouping (46,890).
+P(lower-2) represents the probability associated with the larger lower bound (lower-2) of the one-sided randomized CI.
 7 Conclusion
-This paper develops a simple, statistically rigorous, and computationally efficient framework for assessing feature significance in supervised learning.
+This paper develops a simple and computationally efficient framework for assessing feature significance in supervised learning.
 Our randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation.
 Empirical studies show strong performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, demonstrating its ability to inform high-stakes decisions with legal, economic, and regulatory implications.
-More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines rigor with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
+More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines statistical grounding with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
 Appendix A Proofs of technical results
@@
 The independence of $\epsilon$ and $X$ then implies conditional independence. ∎
+Appendix B Panel data
+Consider panel data $\{Z_{i,t}\}$ with $i=1,\ldots,m$ indexing units and $t=1,\ldots,T$ indexing time.
+We partition $\{(i,t):i\in[m],t\in[T]\}$ into two disjoint subsets, $I_{1}$ and $I_{2}$, such that each pair belongs to exactly one subset.
+This formulation allows for commonly used data-splitting schemes, such as temporal splits (distinguishing between in-sample and out-of-sample time periods), unit splits (assigning a unit’s entire time series to either $I_{1}$ or $I_{2}$), or combinations of both.
+In the case of an unbalanced panel, where units may be observed at different points in time, only observed pairs $(i,t)$ are assigned to $I_{1}$ or $I_{2}$.
+B.1 Conditioning
+Conditioning entails focusing on an appropriate subset $I\subseteq I_{2}$ when constructing the feature effect.
+For example, in the credit card analysis of Section 6.1, we condition on card applicants who are married to understand feature significance for a group of borrowers with certain features.
+If some samples in $I_{2}$ are dependent, as is often the case for panel data, we can choose $I$ as the subset of independent samples $(i,t)\in I_{2}$ (albeit at the cost of reducing the size of the test set).
+We can further partition $I$ by period $t$, to construct feature effects $\{\Delta_{\cdot,t}:(\cdot,t)\in I\}$ indexed by $t$.
+This approach leads to a sequence of tests, and a time series of $p$-values as well as importance scores and their CIs.
+The time-series behavior may provide useful insight.
+For the purpose of constructing baseline and comparison features, the training set $I_{1}$ is also treated as a collection of $|I_{1}|$ samples–with unit and time information tied to each sample–fully harnessing panel data richness.
+For each $(i,t)\in I_{2}$, the baseline features can harness information from not only all training samples $I_{1}$ but also testing samples corresponding to the same unit $\{X_{i,\cdot}:(i,\cdot)\in I_{2}\}$.
+This flexibility enables unit-specific and time-dependent baselines such as lagged feature value (e.g., last month’s feature value) and historical aggregation, enhancing the baseline’s meaningfulness.
+B.2 Grouping
+An alternative grouping approach allows for joint testing across the time-series.
+The idea is to aggregate feature effects within a unit of dependence, so that dependence is contained within groups.
+We treat the testing set $I_{2}$ as a collection of $|I_{2}|$ samples and compute the baseline and treatment predicted responses: $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, $(i,t)\in I_{2}$, where $\underline{X}_{i,t}$ and $\overline{X}_{i,t}$ are baseline and treatment feature values constructed as in the conditional method.
+Note that this differs slightly from the conditional approach.
+In the grouping approach, $I_{2}$ is treated as a collection of $|I_{2}|$ samples only until the construction of the baseline and comparison predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, unlike the conditional approach which includes the construction of the feature effect $\Delta_{i,t}$.
+Then, we group the computed predictions along with the corresponding true responses by unit to compute the unit-level feature effect
+$\Delta_{i}=S(\overline{f}_{i},Y_{i})-S(\underline{f}_{i},Y_{i})$
+where $S$ is a trajectory-level score function, $\underline{f}_{i}=\{f(\underline{X}_{i,t}):(i,t)\in I_{2}\}$, $\overline{f}_{i}=\{f(\overline{X}_{i,t}):(i,t)\in I_{2}\}$, and $Y_{i}=\{Y_{i,t}:(i,t)\in I_{2}\}$.
+Unlike previous single-period settings where the score function $S$ evaluates only one prediction against one true response, the trajectory-level score function evaluates the entire time series of predictions against the time series of true responses.
+The trajectory-level score function can extend single-period score functions.
+An example is the negative mean squared error, where the average is across time steps.
+More complex score functions can harness rich temporal data, like correlation (Information Coefficient, IC), negative error in maximum drawdown (MDD): S(m,y)=|MDD(y)MDD(m)|S(m,y)=-|MDD(y)-MDD(m)|, or other time series evaluations suitable for specific domain usage.
+More broadly, the grouping approach extends beyond temporal dependence structures.
+For instance, for spatially dependent data, one could group testing samples by region or neighborhood, under the assumption that neighborhoods are (approximately) independent.
+Similarly, in settings with repeated measurements, grouping by the underlying entity or context may be effective.
+B.3 Hybrid approach
+Conditioning and grouping can be combined.
+For example, suppose we want to analyze a feature’s impact on model performance over a period of time.
+A long period avoids short-term noise but renders single-period conditioning insufficient.
+A shorter period ensures the relevance of data but may rule out grouping.
+This can be addressed by first conditioning on a time frame of desirable length and then grouping the conditional data by units corresponding to each time-unit pair.
+As with conditioning, this procedure can be repeated across different time windows, leading to a rolling-window analysis with smoothed single-period results.
+Another notable use case of the hybrid method is the analysis of per-unit consistency of feature performance: i.e., for each unit, how consistent across time is a feature’s impact on model performance?
+A longer window reduces short-term fluctuations; when the multi-period score function is an aggregation of some single-period scores, a longer window also introduces another interesting aspect: per-unit consistency.
+For example, suppose we choose the multi-period score function to simply be the average of a single-period function, such as negative mean squared error.
+Then, it can be shown that the unit-level feature effect $\Delta_{i}$ is essentially the average of unit-time-level feature effects $\Delta_{i,t}$ and, therefore, embeds some degree of consistency across time.
+The longer the time window, the higher the degree of consistency imposed, as the bar is set higher.
+Appendix C Missing Data
+In practice, many data sets have missing (feature) values.
+Modelers often impute missing values using a variety of approaches.
+Sometimes missingness indicator features are added to distinguish imputed from actually observed values in the model training process.
+In general, we recommend dropping samples with missing/imputed feature values from the test set, assuming these samples can be identified.
+While this approach reduces the size of the test set and hence the power of the test, it eliminates potentially biased feature effect samples (e.g., feature effects equal to 0, arising when a baseline value is identical to the imputed value such as the mean).
+When applying the grouping approach discussed in Appendix B.2, the removal can be performed at the prediction level $f(\cdot)$ instead of the feature effect level $\Delta_{i}$ to preserve sample size.
+For example, in a panel setting, if a unit $i$ has missing feature values at only one time step tt—i.e., $X^{\ell}_{i,t}$ is missing, but $X^{\ell}_{i,s}$ for $s\neq t$ are present—it may suffice to drop just the predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$ for that specific time step from the group of unit $i$’s predictions.
+In other words, if only a single point is missing from the time series of feature values, only drop the corresponding point from the time series of predictions, not the entire time series.
+However, this situation should not be confused with the case when the data is missing in the baseline value but still present in the original data: i.e., $\underline{X}_{i}$ is missing, but $X^{\ell}_{i}$ is present.
+For example, when using a lagged baseline and data are missing at time t1t-1 but are present at time $t$.
+Additionally, a common practice in handling missing data is to add a missingness flag–an additional dummy variable indicating missing feature values.
+To ensure schematic consistency of the baseline values, the missing flag should always be updated to correspond to current value of the corresponding feature.

### v4

@@
 Abstract
-The opacity of many supervised learning algorithms remains a key challenge, hindering scientific discovery and limiting broader deployment—particularly in high-stakes domains.
-This paper develops model- and distribution-agnostic significance tests to assess the influence of input features in any regression or classification algorithm.
-Our method evaluates a feature’s incremental contribution to model performance by masking its values across samples.
-Under the null hypothesis, the distribution of performance differences across a test set has a non-positive median.
-We construct a uniformly most powerful, randomized sign test for this median, yielding exact $p$-values for assessing feature significance and confidence intervals with exact coverage for estimating population-level feature importance.
-The approach requires minimal assumptions, avoids model retraining or auxiliary models, and remains computationally efficient even for large-scale, high-dimensional settings.
-Experiments on synthetic tasks validate its statistical and computational advantages, and applications to real-world data demonstrate its ability to inform high-stakes decisions with legal, economic, and regulatory implications.
+Machine learning has become a central tool across scientific, industrial, and policy domains.
+Algorithms now identify chemical properties, forecast disease risk, screen borrowers, and guide public interventions.
+Yet this predictive power often comes at the cost of transparency: we rarely know which input features truly drive a model’s predictions.
+Without such understanding, researchers cannot draw reliable scientific conclusions, practitioners cannot ensure fairness or accountability, and policy makers cannot trust or govern model-based decisions.
+Despite its importance, existing tools for assessing feature influence are limited—most lack statistical guarantees, and many require costly retraining or surrogate modeling, making them impractical for large modern models.
+We introduce AICO, a broadly applicable framework that turns model interpretability into an efficient statistical exercise.
+AICO asks, for any trained regression or classification model, whether each feature genuinely improves model performance.
+It does so by masking the feature’s information and measuring the resulting change in performance.
+The method delivers exact, finite-sample inference—exact feature $p$-values and confidence intervals—without any retraining, surrogate modeling, or distributional assumptions, making it feasible for today’s large-scale algorithms.
+In both controlled experiments and real applications—from credit scoring to mortgage-behavior prediction—AICO consistently pinpoints the variables that drive model behavior, providing a fast and reliable path toward transparent and trustworthy machine learning.
 1 Introduction
-Supervised learning algorithms have become foundational tools across scientific and industrial domains, enabling breakthroughs and transforming key sectors.
-However, their opacity poses a major challenge, hindering scientific discovery and limiting adoption in high-stakes domains like finance, public policy and health care, where transparency is crucial.
-This paper proposes novel testing methods for addressing the opacity of learning algorithms, focusing on the role of an algorithms’s input variables (features).
-Specifically, we develop a family of hypothesis tests to assess the significance of features in any supervised learning model, whether for regression or classification.
-The tests are model- and distribution-agnostic: they only require the ability to query the trained model—no assumptions about the model class, data distribution or model target (e.g.
-conditional mean, quantile) are needed.
-The computational efficiency of our tests is unprecedented.
-The procedure does not require retraining or surrogate models, making it well-suited for analyzing complex, high-dimensional algorithms that are costly to train.
-The tests are based on a new approach for evaluating the impact of a feature.
-We quantify a feature’s incremental contribution to a user-defined score function, which measures a model’s predictive or decision-making performance on a test sample.
-Specifically, we compute the feature effect as the change in score when the model is given the actual value of the feature versus when the feature is masked with a baseline value.
-This masking approach, which we call Add-In Covariates (AICO), avoids the variance inflation (excess randomness) introduced by permutation methods and the high cost of retraining-based and surrogate-model-based strategies.
-The score function and masking baseline can be user-specified to align with specific interpretability goals and downstream model use cases; we discuss a number of choices and the associated tradeoffs.
-The feature effect measures the impact of a feature’s variability on model score; it represents the importance of a feature at the level of an individual sample.
-A null-feature’s effect has a distribution across a collection of test samples whose median is non-positive.
-We construct a uniformly most powerful, randomized sign test for this median.
-Randomization at the boundary of the critical region ensures the test’s exactness, i.e., its size (type I error probability) matching the user-specified nominal level (e.g.
-$5\%$).
-The test delivers exact, randomized feature $p$-values that have, conditionally on the data, a uniform distribution on an interval which can be reported in practice.
-It is simple, robust and permits an efficient implementation.
-It requires minimal assumptions, ensuring its broad validity in practice.
-These properties allow users to make valid inferences regarding a model’s features without relying on large-sample approximations or parametric assumptions.
-The median feature effect being tested is a natural measure of feature importance at the population level.
-We develop randomized, exact-coverage confidence intervals for the median that, together with the $p$-values, enable principled inference regarding feature influence.
-The availability of confidence intervals distinguishes the median from widely-used algorithmic measures of feature importance (see Section 1.1 below).
-Moreover, depending on the selection of the masking baseline, the median can have several useful properties.
-These include robustness to correlation between features, which is common in practice.
-Furthermore, for linear regression, the median is proportional to the square of the feature’s regression coefficient, which is the canonical measure of importance in the linear context.
-Finally, we provide conditions guaranteeing that a feature’s median being 0 is equivalent to it being conditionally independent of the response given the remaining features.
-Extensive experiments on synthetic “known truth” tasks demonstrate the statistical and computational performance of our approach in nonlinear regression and classification settings.
-We compare our tests with the recently developed, model-agnostic and non-asymptotic tests of ?, ?
-and ?.
-We also compare with the model-X knockoff procedures of ?
+Machine learning has become integral to modern science, industry, and public policy.
+Algorithms now predict molecular activity, identify disease markers, forecast financial defaults, and guide the allocation of social resources.
+Yet even as these systems achieve extraordinary predictive accuracy, their inner workings remain largely opaque.
+We often do not know which input features truly shape a model’s predictions, which merely correlate by chance, or which may encode unwanted biases.
+This opacity limits scientific understanding, hinders regulatory oversight, and undermines public trust in data-driven decision making.
+Despite extensive progress in explainable artificial intelligence, most tools for measuring feature importance lack statistical guarantees or depend on computationally costly retraining or surrogate modeling procedures, restricting their use in today’s large, complex models.
+As a result, reconciling interpretability with scalable inference remains an open challenge for modern machine-learning practice.
+We introduce AICO, a tractable method for feature significance testing in any supervised learning model, whether for regression or classification.
+AICO transforms the question of interpretability into one of formal statistical inference: it tests whether a feature contributes meaningfully to a model’s predictive or decision-making performance, using only the model’s fitted predictions and observed outcomes.
+AICO constructs test statistics based on the change in model performance when a feature is masked.
+It is model- and distribution-agnostic: it makes no assumptions about the model class, data distribution, or estimand (for example, conditional mean or quantile).
+The method does not rely on retraining or surrogate models, making it well suited for analyzing complex, high-dimensional algorithms that are costly to train.
+It bridges a long-standing gap between machine-learning practice and classical statistical inference, offering a tractable route to transparent, trustworthy prediction in today’s large-scale applications.
+The AICO feature effect—the change in model performance due to masking—quantifies a feature’s importance for each individual sample.
+For a null feature, these effects form a distribution across test samples whose median is non-positive.
+A uniformly most powerful randomized sign test achieves finite-sample exactness: its size (type-I error probability) exactly matches the nominal level (e.g., $5\%$).
+The procedure produces exact, randomized feature $p$-values whose conditional distribution, given the data, is uniform on an interval that can be reported in practice.
+It is simple, robust, and computationally efficient, relying on minimal assumptions and remaining valid under broad conditions.
+These properties enable valid inference on a model’s features without relying on large-sample approximations or parametric assumptions.
+The median feature effect being tested provides a natural, population-level measure of feature importance.
+We develop randomized, exact-coverage confidence intervals for this median that, together with the $p$-values, enable principled inference on a feature’s influence.
+The availability of confidence intervals distinguishes the median from widely used algorithmic measures of feature importance (see Section 1.1).
+The median also has several desirable properties that depend on the choice of masking value.
+It can be robust to correlation among features—a pervasive challenge in practical data settings—and, in linear regression, it is proportional to the square of the feature’s regression coefficient, the canonical measure of importance in the linear context.
+Finally, we provide conditions under which a feature’s median equal to zero is equivalent to conditional independence of that feature and the response given the remaining covariates.
+Extensive experiments on synthetic “known-truth” tasks demonstrate the statistical and computational performance of our approach in nonlinear regression and classification settings.
+We compare AICO with powerful, model-agnostic, and non-asymptotic tests by ?, ?, and ?, as well as with the model-X knockoff procedures of ?
 and ?
-for variable selection with false discovery control.
-Our test is found to be as statistically powerful as the alternatives, with minimal false discovery, while reducing computational costs by several orders of magnitude.
-Moreover, we show that our feature importance measure is robust to noise in the data, delivering stable feature importance rankings.
-This contrasts sharply with widely-used algorithmic feature importance measures including SHAP (?) which are found to be highly sensitive to noise in the data, leading to unstable feature rankings that make it difficult to draw reliable inferences regarding variable importance.
-Our findings complement the discussion of the theoretical caveats of algorithmic importance measures in ?, ?
-and others.
+for variable selection under false discovery control.
+Our test achieves comparable or higher statistical power than these alternatives, with minimal false discoveries, while reducing computational cost by several orders of magnitude.
+We further show that the AICO feature-importance measure is robust to noise in the data, yielding stable feature rankings.
+This stands in sharp contrast to widely used algorithmic importance measures such as SHAP (?), which are highly sensitive to data perturbations and can produce unstable rankings that hinder reliable inference about variable importance.
+These empirical findings reinforce recent theoretical discussions of the limitations of algorithmic importance measures in ?, ?, and related work.
 Empirical analyses demonstrate the ability of our framework to inform high-stakes decisions with legal, economic, and regulatory implications.
-In the context of a credit scoring task using large ensembles of gradient boosting machines, our approach identifies and ranks several significant features describing past payment behavior and social factors.
-Interestingly, we find that gender, marital status, and age of card holders (all of which are protected characteristics under US lending laws) are significant predictors of card default, after controlling for past payment behavior.
-In the context of a mortgage state transition prediction task using a recently proposed set-sequence model based on long-convolution architectures (?) and trained on extensive panel data, our test identifies time-varying macro-economic variables such as mortgage rates, as well as features summarizing past state transitions as significant predictors of borrower behavior.
-The rest of this introduction discusses the related literature.
-Section 2 introduces the notion of feature effects while Section 3 develops the testing procedure.
+In a credit scoring task using ensembles of gradient-boosting models, AICO identifies and ranks several statistically significant features describing past payment behavior and social factors.
+Notably, the analysis reveals that gender, marital status, and age of card holders—protected characteristics under U.S.
+lending laws—are statistically significant predictors of card default, even after controlling for past payment behavior.
+In a mortgage state–transition prediction task based on a recently proposed set–sequence model with long-convolution architectures (?) and trained on extensive panel data, AICO detects time-varying macroeconomic variables such as mortgage rates, as well as features summarizing past state transitions, as significant predictors of borrower behavior.
+The remainder of this introduction discusses the related literature.
+Section 2 introduces the notion of feature effects, and Section 3 develops the testing procedure.
 Section 4 analyzes the median feature effect as an importance measure dual to the test.
-Section 5 studies the performance of the test and importance measure in synthetic regression and classification settings.
-Section 6 deploys the test on two real-world data sets.
+Section 5 examines the performance of the test and importance measure in synthetic regression and classification settings, while Section 6 applies the method to two real-world data sets.
 Section 7 concludes.
-The proofs of technical results are given in an appendix.
-A Python package implementing our testing and importance procedures is available at https://github.com/ji-chartsiri/AICO.
+Proofs of the theoretical results are provided in an Appendix.
+An open-source Python package implementing the testing and importance procedures is available at https://github.com/ji-chartsiri/AICO.
 1.1 Related literature
 There is a large body of work developing algorithmic feature importance measures; see ?
-for an overview.
+for an overview and ?
+for a unifying framework for interpretability methods.
 Popular examples include measures based on Shapley values (e.g.
@@
 We demonstrate that algorithmic measures can be highly susceptible to noise in the data so the lack of confidence intervals is concerning.
-Moreover, algorithmic measures such as SHAP can be biased in the presence of feature correlation (?).
+Moreover, algorithmic measures are often biased in the presence of feature correlation (?, ?).
 Our approach can deliver importance measures that are robust to feature correlation.
 Finally, we demonstrate that our method achieves orders-of-magnitude improvements in computational efficiency over widely used algorithmic importance approaches, such as SHAP and permutation scores.
-The statistical and econometric literature has developed significance tests for the variables in nonparametric regression models.
-?, ?, ?, ?, ?
-and others employ kernel methods to construct asymptotic-based testing procedures.
-?, ?
-and ?
-and others construct gradient-based asymptotic tests using neural networks.
-Unlike these tests, ours are model-agnostic, non-asymptotic, deliver exact $p$-values and confidence intervals, and can accommodate a wide range of data structures.
-Pursuing a model-agnostic approach, ?, ?, ?, ?, ?, ?
-and others develop non-asymptotic tests for a conditional independence hypothesis of feature significance in the model-X knockoff setting of ?.
-[^2] Conditional independence testing is challenging: ?
-show that no test of conditional independence can be uniformly valid against arbitrary alternatives.
-We show that, for an appropriate choice of the masking baseline, our null is generally weaker than the conditional independence null studied in these papers; equivalence holds under additional conditions.
+The statistical and econometric literature has long studied significance testing for variables in nonparametric regression models.
+?, ?, ?, ?, ?, and others employ kernel-based methods to construct asymptotic testing procedures.
+?, ?, ?, and others develop gradient-based asymptotic tests using neural networks.
 ?
-construct non-asymptotic tests based on the drop in model predictive accuracy due to removing features (LOCO: Leave-out-covariates).
-We, in contrast, test the improvement in model score due to masking a feature.
-The computational requirements of the aforementioned model-agnostic testing procedures are a limiting factor.
-The LOCO test requires re-training the model for each feature to be tested, which can be prohibitive for feature-rich models that are costly to train.
-The model-X knockoff procedures require constructing an auxiliary model of the joint feature distribution, which can be as challenging and computationally costly as constructing the model to be tested.
-In fact, ?
+develop permutation-based asymptotic tests for random forests; see also ?
+for an alternative approach using $U$-statistics.
+In contrast, the AICO tests are model-agnostic and non-asymptotic, deliver exact $p$-values and confidence intervals, and accommodate a broad range of data structures.
+Pursuing a model-agnostic approach, ?, ?, ?, ?, ?, ?, and others develop non-asymptotic tests for conditional independence in the model-X knockoff framework of ?.
+Conditional independence testing is inherently difficult: ?
+show that no such test can be uniformly valid against arbitrary alternatives.
+We show that, for an appropriate choice of the feature masking value, the AICO null is generally weaker than the conditional-independence null studied in these papers, with equivalence holding under additional conditions.
+?
+suggest model-agnostic tests based on the drop in model predictive accuracy when features are removed from the model (LOCO: leave-out-covariates; see also ?
+and ?).
+In contrast, AICO tests the change in model performance due to masking a feature.
+The computational demands of the aforementioned model-agnostic testing procedures are a major limiting factor.
+The LOCO test requires retraining the model for each feature to be tested, which can be prohibitive for feature-rich models that are costly to train.
+The model-X knockoff procedures require constructing an auxiliary model of the joint feature distribution, a task that can be as challenging and computationally expensive as fitting the predictive model itself.
+Indeed, ?
 show that no algorithm can efficiently compute nontrivial knockoffs for arbitrary feature distributions.
-Our approach does not require any model re-training or the construction of auxiliary models, delivering an unprecedented degree of computational efficiency.
-Our numerical results demonstrate that the drastic gains in computational performance we achieve do not come at the expense of statistical performance.
+In contrast, AICO requires neither model retraining nor auxiliary modeling, delivering an unprecedented degree of computational efficiency.
+Our numerical experiments demonstrate that these drastic gains in computational speed do not come at the expense of statistical power or inferential accuracy.
 2 Feature effect
-We consider data $Z_{1},\ldots,Z_{m}$ with $Z_{i}=(X_{i},Y_{i})$.
-Here, $X_{i}=(X^{1}_{i},\ldots,X^{d}_{i})$ is a $d$-dimensional feature vector taking values in a set $\mathcal{X}$ and $Y_{i}$ is a response taking values in a set $\mathcal{Y}$.
-In a typical regression setting with tabular data, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$.
+Consider data $Z_{1},\ldots,Z_{m}$ with $Z_{i}=(X_{i},Y_{i})$.
+Here, $X_{i}=(X^{1}_{i},\ldots,X^{d}_{i})$ is a feature vector taking values in a set $\mathcal{X}$, and $Y_{i}$ is a response taking values in a set $\mathcal{Y}$.
+For example, in a typical regression setting with tabular data, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$.
 We make no assumptions on the distribution of the $Z_{i}$’s.
-We take as given a random split of $m]=\{1,\ldots,m\}$ into non-intersecting subsets $I_{1}$ and $I_{2}$, which are used for training/validation and testing purposes, respectively.
-This partitioning is standard in supervised learning settings.
-In our testing framework, data splitting facilitates post-selection inference.
+We take as given a random split of $m]=\{1,\ldots,m\}$ into non-intersecting subsets $I_{1}$ and $I_{2}$, used for training and testing, respectively.
+Such data splitting is standard in supervised learning and, in our setting, facilitates post-selection inference.
 We also take as given a model (algorithm) $f:\mathcal{X}\to\mathcal{M}$ trained on $\{Z_{i}\}_{i\in I_{1}}$, which is to be tested.
-We do not make any assumptions on the form of the modeling target; $f$ could represent $\mathbb{E}(Y_{i}|X_{i})$, a conditional quantile, or some other statistic of $Y_{i}$.
-We also do not impose assumptions on the training criterion (eg.
-$L_{2}$ error) or approach (eg.
-empirical risk minimization) being used for training.
-We would like to test the statistical significance of a feature $\ell\in[d]$ under the model $f$.
-(Unless noted otherwise, $\ell$ is fixed.) In the special case that $f$ is linear, standard asymptotic procedures such as the $t$-test are available.
-We develop a simple, non-asymptotic and model-agnostic testing procedure that does not require a special model structure, and is valid much more broadly for models $f$ of essentially any form.
-We consider the impact on model performance of a feature’s variability across the test set $I_{2}$.
-Performance is measured by a deterministic score function $S:\mathcal{M}\times\mathcal{Y}\to\mathbb{R}$, which relates model output to response.
-Specifically, for a given feature value $x\in\mathcal{X}$ and response value $y\in\mathcal{Y}$, the score $S(f(x),y)$ evaluates the quality of the model output $f(x)$ in view of $y$.
-Higher scores reflect better performance.
-A natural choice for $S$ is the negative of the loss function used for the training of $f$.
-Examples include the negative $L_{2}$ loss $S(m,y)=-(m-y)^{2}$ in a regression context, the negative pinball loss in a quantile regression context, or the negative (binary or multi-class) cross-entropy loss in a classification context.
-With these specifications $S$ measures predictive accuracy.
-Other specifications of $S$ can measure model performance (e.g.
-revenue) in the context of a downstream screening or decisioning application of $f$, for example the financial loss caused by a poor decision due to mis-classification.
-Finally, we can specify a function $S$ that does not depend on the second argument representing the response, to measure the behavior (e.g.
-sensitivity) of the model output on its own, without reference to the response.
-We evaluate the effect on model score of a feature’s variability across the test set.
-For each $i\in I_{2}$ we construct from $X_{i}$ a baseline feature value $\underline{X}_{i}\in\mathcal{X}$ serving as a reference and a comparison feature value $\overline{X}_{i}\in\mathcal{X}$.
-In the baseline, the value of the feature to be tested is masked with a reference value whose choice is further discussed below.
-The comparison value $\overline{X}_{i}$ is unmasked so that the predicted response $f(\overline{X}_{i})$ can harness the actual sample $i$ value of the feature to be tested.
-We propose to measure the feature effect by the resulting score improvement.
+We make no assumptions on the form of the estimand; $f$ may represent the conditional mean, a quantile, or another statistic of $Y_{i}$.
+We likewise do not impose assumptions on the training criterion (e.g., $L_{2}$ error) or on the optimization approach (e.g., empirical risk minimization) used to fit $f$.
+We aim to test the statistical significance of a feature $\ell\in[d]$ under the fitted model $f$.
+(Unless noted otherwise, $\ell$ is fixed.) When $f$ is linear, classical asymptotic procedures such as the $t$-test are available.
+We develop instead a simple, non-asymptotic, and model-agnostic testing procedure that does not rely on a specific model structure and remains valid for models $f$ of essentially any form.
+We consider the impact of a feature’s variability on model performance across the test set $I_{2}$.
+Performance is measured by a deterministic score function $S:\mathcal{M}\times\mathcal{Y}\to\mathbb{R}$ that relates model output to response.
+For a given feature value $x\in\mathcal{X}$ and response value $y\in\mathcal{Y}$, the score $S(f(x),y)$ evaluates the quality of the model output $f(x)$ relative to $y$, with higher scores indicating better performance.
+A natural choice for $S$ is the negative of the loss function used to train $f$.
+Examples include the negative $L_{2}$ loss $S(m,y)=-(m-y)^{2}$ in regression, the negative pinball loss in quantile regression, or the negative (binary or multi-class) cross-entropy loss in classification.
+With these specifications, $S$ measures predictive accuracy.
+Other choices of $S$ can quantify model performance in downstream screening or decisioning applications of $f$ (e.g., revenue or cost), such as the financial loss caused by a poor decision due to misclassification.
+Finally, one may specify a function $S$ that does not depend on the second argument $y$ to measure the behavior of the model output itself (for example, its sensitivity to the inputs).
+We evaluate the effect of a feature’s variability on the model score across the test set.
+For each $i\in I_{2}$, we construct from $X_{i}$ a baseline feature value $\underline{X}_{i}\in\mathcal{X}$, serving as a reference, and a comparison feature value $\overline{X}_{i}\in\mathcal{X}$.
+In the baseline, the value of the feature to be tested is masked with a reference value, the choice of which is discussed further below.
+The comparison value $\overline{X}_{i}$ is unmasked so that the predicted response $f(\overline{X}_{i})$ can leverage the actual sample-$i$ value of the feature being tested.
+We measure the feature’s effect by the corresponding improvement in model score.
 Definition 2.1 (Feature Effect).
-For a given score function $S$, baseline feature value $\underline{X}_{i}$ and comparison feature value $\overline{X}_{i}$, the feature effect on test sample $i\in I_{2}$ is the score delta
+For a given score function $S$, baseline feature value $\underline{X}_{i}$, and comparison feature value $\overline{X}_{i}$, the feature effect on test sample $i\in I_{2}$ is defined as the score difference
 $\Delta_{i}=S(f(\overline{X}_{i}),Y_{i})-S(f(\underline{X}_{i}),Y_{i}).$
-Note that masking feature values for purposes of measuring a feature effect has significant advantages over alternative approaches such as randomly shuffling feature values (e.g.
-?
-and ?) or dropping a feature from the feature vector.
-Shuffling introduces excess variance into the measurement procedure and dropping features requires training sub-models, which (drastically) reduces the computational efficiency of the measurement procedure.
-Our masking approach avoids these issues.
+Masking feature values to measure a feature’s effect offers significant advantages over alternatives such as random shuffling, random perturbation or dropping the feature from the vector.
+Random procedures introduce additional variance into the measurement, while dropping features requires training sub-models, which drastically reduces computational efficiency.
+Our masking approach avoids both issues, yielding stable and computationally efficient estimates of feature importance.
 2.1 Baseline and comparison features
-Along with the selection of the score function $S$, the selection of baseline and comparison feature values enables users to tailor the feature effect to the model and data structures at hand as well as the testing objectives.
-We provide a conceptional framework and several examples.
-Throughout, fix a feature $\ell\in[d]$ that is to be tested; our notation usually suppresses the dependence on $\ell$.
-We start by selecting a non-empty set $\underline{I}\subseteq I_{1}$ of training samples that will be used to construct the baseline $\underline{X}_{i}$ for the test sample $i\in I_{2}$.
+Along with the selection of the score function $S$, the choice of baseline and comparison feature values allows users to tailor the feature effect to the model, data structure, and testing objectives at hand.
+We provide a conceptual framework and several examples.
+Throughout, fix a feature $\ell\in[d]$ to be tested; our notation usually suppresses the dependence on $\ell$.
+We begin by selecting a non-empty set $\underline{I}\subseteq I_{1}$ of training samples that will be used to construct the baseline $\underline{X}_{i}$ for each test sample $i\in I_{2}$.
 In a second step, we aggregate the samples in $\underline{I}$ to obtain $\underline{X}_{i}$.
-Alternative aggregators include the mean, a quantile or a mode of $\{X_{k}:k\in\underline{I}\}$.
+Possible aggregation rules include the mean, a quantile, or a mode of $\{X_{k}:k\in\underline{I}\}$.
 Example 2.1.
-Take $\underline{I}=I_{1}$ and let $\underline{X}_{i}=(v^{1},\ldots,v^{d})$ for all $i\in I_{2}$ where $v^{j}$ is the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ for each $j\in[d]$.
-The predicted response $f(\underline{X}_{i})$ is constant across the test set.
-If feature $\ell=1$ is to be tested, select the comparison value $\overline{X}_{i}=(X_{i}^{1},v^{2},\ldots,v^{d})$.
-The predicted response $f(\overline{X}_{i})$ harnesses feature 1’s actual sample value, $X_{i}^{1}$.
-The feature effect $\Delta_{i}$ represents the associated score improvement.
-An alternative choice is $\underline{X}_{i}=(v^{1},X^{2}_{i},\ldots,X^{d}_{i})$ and $\overline{X}_{i}=X_{i}=(X_{i}^{1},\ldots,X^{d}_{i})$, again supposing feature $\ell=1$ is to be tested.
-The predicted baseline response $f(\underline{X}_{i})$ is no longer constant across the test set, and accounts for the variation of features $X^{2}_{i},\ldots,X^{d}_{i}$.
-Here $\Delta_{i}$ measures a conditional feature effect.
-Choices other than the mean may be appropriate for $v^{j}$.
-For example, in the case of a one-hot feature $X^{j}_{i}\in\{0,1\}$ which might encode a categorical variable, sensible choices for $v^{j}$ are a mode of $\{X^{j}_{k}:k\in\underline{I}\}$ or the value $1-X^{j}_{i}$.
-[^3] See ?
-for extensions.
-Consider choices of $\underline{I}$ beyond the entire training set $I_{1}$.
-We can pick specific samples from $I_{1}$ by first fixing a reference point $r$ in $\mathcal{Y}$, $\mathcal{M}$, or $\mathcal{X}$ and then selecting the samples from $I_{1}$ that lie in or outside a specified neighborhood of $r$.
-The point $r$ can represent the (weighted) mean, mode or a $\lambda$-quantile of the responses $\{Y_{k}:k\in I_{1}\}$, the predicted responses $\{f(X_{k}):k\in I_{1}\}$, or the values of the feature $\ell$ to be tested, $\{X^{\ell}_{k}:k\in I_{1}\}$.
-Potentially interesting special cases are quantiles for $\lambda\in\{0,1\}$.
-Concretely, let $Y_{r}\in\mathcal{Y}$ be a reference response value and $d$ be a distance measure on $\mathcal{Y}$.
-We can now take $\underline{I}$ as the set of samples $k\in I_{1}$ for which $d(Y_{k},Y_{r})$ does not exceed or surpasses some threshold value.
-Other criteria can be envisioned, for example samples $k\in I_{1}$ for which $Y_{k}$ exceeds or falls short of $Y_{r}$.
-A similar approach applies to reference points in $\mathcal{M}$ or $\mathcal{X}$.
+Take $\underline{I}=I_{1}$ and let $\underline{X}_{i}=(v^{1},\ldots,v^{d})$ for all $i\in I_{2}$, where $v^{j}$ is the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ for each $j\in[d]$.
+Then $f(\underline{X}_{i})$ is constant across the test set.
+For $\ell=1$, take $\overline{X}_{i}=(X_{i}^{1},v^{2},\ldots,v^{d})$.
+Here, the feature’s actual sample value $X_{i}^{1}$ is added back in while all other components remain fixed—an operation that gives rise to the name AICO (Add-In COvariates).
+The predicted response $f(\overline{X}_{i})$ therefore leverages the true variation in feature 1, and the feature effect $\Delta_{i}$ represents the resulting improvement in model score.
+An alternative choice is $\underline{X}_{i}=(v^{1},X_{i}^{2},\ldots,X_{i}^{d})$ and $\overline{X}_{i}=X_{i}=(X_{i}^{1},\ldots,X_{i}^{d})$, again with $\ell=1$.
+Here, the predicted baseline response $f(\underline{X}_{i})$ is no longer constant across the test set, as it incorporates variation in the remaining features $X_{i}^{2},\ldots,X_{i}^{d}$.
+Adding the true value $X_{i}^{1}$ back into this context performs the same AICO operation, now yielding a conditional feature effect $\Delta_{i}$ that measures the incremental contribution of feature 1 given the others.
+Choices other than the mean may also be appropriate for $v^{j}$.
+For instance, when $X_{i}^{j}\in\{0,1\}$ encodes a categorical variable, the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ may not represent a valid feature value.
+Sensible alternatives include the mode of $\{X^{j}_{k}:k\in\underline{I}\}$ or the complementary value $1-X_{i}^{j}$.
+The set $\underline{I}$ need not equal the entire training set $I_{1}$.
+We can select specific samples from $I_{1}$ by first fixing a reference point $r$ in $\mathcal{Y}$, $\mathcal{M}$, or $\mathcal{X}$ and then choosing the samples from $I_{1}$ that lie in or outside a neighborhood of $r$.
+The point $r$ can represent, for example, the (weighted) mean, mode, or a $\lambda$-quantile of the responses $\{Y_{k}:k\in I_{1}\}$, of the predicted responses $\{f(X_{k}):k\in I_{1}\}$, or of the feature values $\{X^{\ell}_{k}:k\in I_{1}\}$ for the feature being tested.
+Quantiles with $\lambda\in\{0,1\}$ are particularly interesting special cases.
+Concretely, let $Y_{r}\in\mathcal{Y}$ be a reference response value and $d$ a distance measure on $\mathcal{Y}$.
+We can then take $\underline{I}$ as the set of samples $k\in I_{1}$ for which $d(Y_{k},Y_{r})$ does not exceed (or surpasses) a given threshold.
+Other criteria can also be used, such as selecting samples for which $Y_{k}$ exceeds or falls short of $Y_{r}$.
+A similar approach applies when $r$ is a reference point in $\mathcal{M}$ or $\mathcal{X}$.
 Example 2.2.
-Consider the reference point $\min_{k\in I_{1}}f(Y_{k})$ in $\mathcal{M}$, which is a quantile of the predicted responses $\{f(Y_{k}):k\in I_{1}\}$ at level $\lambda=0$.
-For example, in a binary soft classification setting such as fraud detection or credit scoring where $Y_{i}=0$ represents the preferred (beneficial) outcome, this reference point represents the training samples with minimum predicted risk.
+Consider the reference point $\min_{k\in I_{1}}f(Y_{k})$ in $\mathcal{M}$, which is the $\lambda=0$ quantile of the predicted responses $\{f(Y_{k}):k\in I_{1}\}$.
+In a binary soft-classification setting such as fraud detection or credit scoring, where $Y_{i}=0$ denotes the preferred (beneficial) outcome, this reference point corresponds to the training samples with minimum predicted risk.
 [^2] See ?.
 We can take $\underline{I}=\{\mbox{argmin}_{k\in I_{1}}f(Y_{k})\}$ as the set of these samples and define $\underline{X}_{i}$ and $\overline{X}_{i}$ as in Example 2.1.
-The associated feature effect $\Delta_{i}$ represents the score delta relative to the minimum risk baseline.
-It measures how much excess risk a feature explains.
-A few basic principles guide the selection of reference points and the construction of the set $\underline{I}\subseteq I_{1}$.
-We wish to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, to avoid feature effects $\Delta_{i}$ that are biased low which result in significance tests with low power.555?
-discuss these issues in the context of image classification.
-Similarly, we want to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$.
-Finally, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space.
-It is important to select an appropriate reference point and rule used to aggregate across $\underline{I}$ (see Example 2.1).
-Alternatively, we might select an appropriate baseline value from the set $\{X_{k}:k\in I_{1}\}$, namely an element that is near the raw off-manifold value emerging from the baseline selection procedure.
+The resulting feature effect $\Delta_{i}$ represents the score delta relative to the minimum-risk baseline and measures how much excess risk a feature explains.
+A few general principles guide the selection of reference points and the construction of $\underline{I}\subseteq I_{1}$.
+We aim to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, avoiding feature effects $\Delta_{i}$ that are biased toward zero and lead to significance tests with low power.333?
+discuss related issues in the context of image classification.
+At the same time, we seek to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$.
+Further, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space.
+It is important to choose an appropriate reference point and aggregation rule for $\underline{I}$ (see Example 2.1).
+Alternatively, one may select a suitable baseline value directly from the set $\{X_{k}:k\in I_{1}\}$—namely, an element near the raw off-manifold value produced by the baseline selection procedure.
 [^4] We also allow for the selection and aggregation of alternative baselines for a given sample.
-A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
+A general approach is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
 2.2 Multiple features, panel data and missing data
-We have focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$.
-The approach naturally extends to testing multiple features, i.e.
-subsets of $[d]$, which facilitates the exploration of higher-order effects including feature interactions.
-Panel data settings, where a collection of units are observed over some time period, are prevalent in many domains.
-Appendix B discusses two simple extensions—conditioning and grouping—to construct the feature effect for panel data.
-Real-world data sets often have missing feature values. Appendix C shows how to treat missing or imputed values when constructing $\Delta_{i}$.
+The discussion above has focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$.
+The same framework naturally extends to subsets of $[d]$, enabling the joint testing of multiple features and the exploration of higher-order effects, including feature interactions.
+Panel data settings—where collections of units are observed over time—are common in many domains.
+Appendix B discusses two simple extensions, conditioning and grouping, that adapt the construction of the feature effect to such data structures.
+Real-world data sets often contain missing feature values. Appendix C describes how missing or imputed values can be handled when constructing $\Delta_{i}$.
 3 Hypothesis testing
-To construct a hypothesis test, we consider the empirical distribution of the feature effect $\Delta_{i}$ across the test samples $i\in I_{2}$ (the feature $\ell\in[d]$ to be tested remains fixed).
-If that distribution has a non-positive center, then the feature is deemed insignificant.
-A large positive center is evidence of a feature being significant.
-There is a battery of alternative tests for the mean, median and other statistics, including $t$, sign, Wicoxon, and permutation tests.
-We propose a simple sign test, which requires minimal assumptions on the data yet offers meaningful asymptotic relative efficiency vs.
-a $t$-test.
+To construct a hypothesis test, we examine the empirical distribution of the feature effects $\Delta_{i}$ across the test samples $i\in I_{2}$, keeping the feature $\ell\in[d]$ fixed.
+If this distribution has a non-positive center, the feature is deemed insignificant; a large positive center provides evidence of significance.
+Several classical tests could be applied to assess such central tendencies, including $t$, sign, Wilcoxon, and permutation tests.
+We propose a simple sign test, which requires minimal assumptions on the data while retaining meaningful asymptotic relative efficiency compared with a $t$-test.
 Assumption 3.1.
-The feature effects $\{\Delta_{i}:i\in I_{2}\}$ are independent draws from a conditional distribution $\mathbb{P}_{M}$ given the training data $I_{1}$ with unique median $M\in\mathbb{R}$.
+The feature effects $\{\Delta_{i}:i\in I_{2}\}$ are independent draws from a conditional distribution $\mathbb{P}_{M}$ given the training data $I_{1}$, with unique median $M\in\mathbb{R}$.
 The associated distribution function $F_{\Delta}$ is continuous.
@@
 Let $B_{n,p}$ be a Binomial-$(n,p)$ variable and let $q_{\lambda}(B_{n,p})=\inf\{x\geq 0:P(B_{n,p}\leq x)\geq\lambda\}$ be its (lower) quantile at $\lambda\in(0,1)$.
-Let $N=|I_{2}|$ be the test sample size.
-Moreover, let $\Pi_{N}(n)=\sum_{i=0}^{n}\pi_{N}(i)$ be the CDF of $B_{N,1/2}$, where $\pi_{N}(i)={N\choose i}2^{-N}$ for $i=0,1,\ldots,N$.
+Let $N=|I_{2}|$ be the size of the test sample, and write $\Pi_{N}(n)=\sum_{i=0}^{n}\pi_{N}(i)$ for the CDF of $B_{N,1/2}$, where $\pi_{N}(i)={N\choose i}2^{-N}$ for $i=0,1,\ldots,N$.
 Proposition 3.2.
@@
 For brevity, we often suppress the dependence of the test statistic $n_{+}$ on $M_{0}$.
-The test rejects $H_{0}$ with probability one if $n_{+}$, the number of feature effects in the test set larger than $M_{0}$, exceeds the $(1-\alpha)$-quantile of the Binomial-$(N,1/2)$ distribution.
-The test randomizes in case $n_{+}$ hits the boundary of the rejection region.
-Relative to a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$, randomization renders the test more aggressive; it ensures maximum power and guarantees exactness in the sense that
+The procedure rejects $H_{0}$ with probability one if $n_{+}$—the number of feature effects in the test set exceeding $M_{0}$—is greater than the $(1-\alpha)$-quantile of the $\mathrm{Binomial}(N,1/2)$ distribution.
+If $n_{+}$ lies exactly on the boundary of the rejection region, the test randomizes.
+Relative to a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$, this randomization makes the procedure slightly more aggressive: it achieves the maximum possible power while guaranteeing exactness, in the sense that
 $\mathbb{E}_{M_{0}}(\phi(n_{+},\alpha))=\alpha,\quad M_{0}\in\mathbb{R}.$ (4)
-Algorithm 1 details the implementation of the test. Our applications of the test in Sections 5 and 6 below will use the value $M_{0}=0$ in (1).
+Algorithm 1 summarizes the implementation of the test. In our empirical applications in Sections 5 and 6, we use the value $M_{0}=0$ in (1).
 In the context of our randomized test the notion of $p$-value is ambiguous.
@@
 4 Measuring feature importance
-We require measures of feature importance that complement the feature $p$-values delivered by our test.
-A local measure of feature importance is given by the feature effect $\Delta_{i}$, which represents the impact of a feature at the sample level.
-The median $M$ of the (conditional) feature effect distribution $F_{\Delta}$ is a natural measure of feature importance at the population level.
+We seek measures of feature importance that complement the feature $p$-values delivered by our test.
+A local measure of feature importance is given by the feature effect $\Delta_{i}$, which captures the impact of a feature at the sample level.
+At the population level, a natural measure of feature importance is the median $M$ of the (conditional) feature-effect distribution $F_{\Delta}$.
 Under Assumption 3.1, a consistent estimator of $M$ is the empirical median
-$\widehat{M}=\mbox{Median}\{\Delta_{i}:i\in I_{2}\},$ (6)
-which satisfies $\mathbb{P}(\widehat{M}<M)=\mathbb{P}(\widehat{M}>M)$ for odd $N=|I_{2}|$ and approximate equality holding for even $N$.
-Thus (6) is (approximately) as likely to overestimate $M$ as to underestimate it.
+$\widehat{M}=\mathrm{Median}\{\Delta_{i}:i\in I_{2}\},$ (6)
+which satisfies $\mathbb{P}(\widehat{M}<M)=\mathbb{P}(\widehat{M}>M)$ for odd $N=|I_{2}|$, with approximate equality for even $N$.
+Hence, (6) is (approximately) as likely to overestimate $M$ as to underestimate it.
 Moreover, the distribution of $\widehat{M}$ is continuous.
@@
 and ?), empirical likelihood (?), sectioning (?) and other methods.
-We develop several useful properties of the population-level feature importance measure defined by the median $M$.
-To this end, let $Z=(X,Y)\sim\mathbb{P}$ where $\mathbb{P}$ denotes the population law from which the observed data points $Z_{i}$ are independently drawn.
-For concreteness, consider a model for estimating the conditional mean, and take $\mu(X)=\mathbb{E}(Y|X)$ and the score function $S$ as given.
-The (population) feature effect is given by $\Delta=S(\mu(\overline{X}),Y)-S(\mu(\underline{X}),Y)$ where $\underline{X}$ and $\overline{X}$ are suitable baseline and comparison features, respectively.
-As discussed in Section 2.1, the selection of $\underline{X}$ and $\overline{X}$ offers some degrees of freedom. Consider the basic choice from Example 2.1,
+We now develop several useful properties of the population-level feature-importance measure defined by the median $M$.
+Let $Z=(X,Y)\sim\mathbb{P}$, where $\mathbb{P}$ denotes the population law from which the observations $Z_{i}$ are independently drawn.
+For concreteness, consider a model that estimates the conditional mean, and let $\mu(X)=\mathbb{E}(Y|X)$ with score function $S$ as given.
+The (population) feature effect is then $\Delta=S(\mu(\overline{X}),Y)-S(\mu(\underline{X}),Y)$, where $\underline{X}$ and $\overline{X}$ denote appropriate baseline and comparison features, respectively.
+As discussed in Section 2.1, the selection of $\underline{X}$ and $\overline{X}$ offers flexibility. Consider the basic choice from Example 2.1,
 $\underline{X}=\mathbb{E}(X)\quad\text{and}\quad\overline{X}=(X^{1},\mathbb{E}(X^{2}),\ldots,\mathbb{E}(X^{d})),$ (10)
-where for convenience we assume, without loss of generality, feature $X^{1}$ is to be tested.
-First note that for the choice (10) and given $\mu$ and $S$, the feature effect $\Delta$ is completely determined by the marginal laws of the features and the joint law of the response $Y$ and $X^{1}$.
-Thus the feature dependence structure does not matter; the importance measure $M$ generated by (10) is robust to feature correlation.
-This is a remarkable property given the sensitivity of popular importance measures including SHAP and LOCO to correlation among features (e.g.
-?).
-The correlation sensitivity induces a bias in these measures, which can lead to dubious importance rankings.
-Next we explore the relationship between $M$ and conditional independence.
-?, ?, ?, ?, ?
-and others develop tests for a conditional independence hypothesis of feature significance in the model-X knockoff setting of ?.
-We show that the hypothesis we test ($M\leq M_{0}=0$) is generally weaker than an alternative conditional independence hypothesis.
-Below, we use the notation $X^{-1}=(X^{2},\ldots,X^{d})$.
+where, without loss of generality, feature $X^{1}$ is to be tested.
+Under this specification (10) and given $\mu$ and $S$, the feature effect $\Delta$ is fully determined by the marginal distributions of the features and the joint distribution of the response $Y$ and feature $X^{1}$.
+Hence, the dependence structure among features does not affect $\Delta$: the importance measure $M$ generated by (10) is robust to feature correlation.
+This robustness is notable given the well-documented sensitivity of popular importance measures—including SHAP and LOCO—to feature correlation (e.g., ?, ?).
+Such correlation sensitivity can introduce bias and lead to misleading importance rankings.
+Next, we explore the relationship between $M$ and conditional independence.
+?, ?, ?, ?, ?, and others develop tests for a conditional-independence hypothesis of feature significance within the model-X knockoff framework of ?.
+We show that the hypothesis tested by AICO, $M\leq M_{0}=0$, is generally weaker than the alternative conditional-independence hypothesis.
+Throughout, we use the notation $X^{-1}=(X^{2},\ldots,X^{d})$.
 Proposition 4.2.
@@
 […]
-$Y=\mu(X)+\epsilon,\;\text{where}\;\mu(X)=\beta\cdot X\;\text{and}\;\beta=(\beta^{1},...,\beta^{d})\in\mathbb{R}^{d}.$ (11)
-Before stating a converse of Proposition 4.2 under (11), we note the following “linear agreement” property: under (11) (and symmetry of the distribution of $\Delta$), the median $M$ measuring feature importance is proportional to the square of the regression coefficient, which is the canonical measure of feature importance in a linear model.
-The proportionality factor is given by the variance of the feature.
+A special case where the converse does hold is the standard linear model
+$Y=\mu(X)+\epsilon,\quad\text{where}\quad\mu(X)=\beta\cdot X,\quad\beta=(\beta^{1},\ldots,\beta^{d})\in\mathbb{R}^{d}.$ (11)
+Before stating the converse of Proposition 4.2 under (11), we note the following “linear agreement” property: under (11) and assuming symmetry of the distribution of $\Delta$, the median $M$ measuring feature importance is proportional to the square of the corresponding regression coefficient.
+This is precisely the canonical measure of feature importance in a linear model, with proportionality factor given by the feature’s variance.
 Proposition 4.3.
@@
 5 Synthetic tasks
-This section illustrates the robust performance of our test and feature importance measure on synthetic “known truth” regression and classification tasks using neural networks.
-It also benchmarks performance against existing model-agnostic tests and variable selection procedures, as well as alternative model-specific and -agnostic feature importance measures.
+This section illustrates the robust performance of our test and feature-importance measure on synthetic “known-truth” regression and classification tasks implemented with neural networks.
+It benchmarks AICO against existing model-agnostic significance tests and variable-selection procedures, as well as against alternative model-specific and model-agnostic measures of feature importance.
 5.1 Regression
-We consider a feature vector $X=(X^{1},\ldots,X^{d})\subseteq\mathbb{R}^{d}$ of dimension $d=19$ and a response $Y\in\mathbb{R}$ of the form
+We consider a feature vector $X=(X^{1},\ldots,X^{d})\subseteq\mathbb{R}^{d}$ of dimension $d=19$ and a response $Y\in\mathbb{R}$ defined by
 $Y=\mu(X)+\epsilon,\quad\epsilon\sim N(0,1),$ (12)
-where the error $\epsilon$ is independent of $X$ and the conditional mean function $\mu$ is given by
+where the error $\epsilon$ is independent of $X$, and the conditional mean function $\mu$ is given by
 $\mu(x)=3$ $+4x_{1}+x_{1}x_{2}+3x_{3}^{2}+2x_{4}x_{5}+6x_{6}+2\sin(x_{7})$
 $+\exp(x_{8})+5x_{9}+3x_{10}+4x_{11}+5x_{12}.$ (13)
-The features $X_{1}$ and $X_{6}$ are jointly normal with mean zero, unit variances and covariances equal to $0.85$.
-The features $X_{2},\ldots,X_{5},X_{7}$ are independent $N(0,1)$, $X_{8}$ is uniform on [1,1][-1,1], $X_{9}=1\{X_{2}+Z<0\}$ where $Z\sim N(0,1)$, $X_{10}$ is Poisson with mean 3 and $X_{11}$ and $X_{12}$ are independent variables with a $t$-distribution and 5 degrees of freedom.
-The non-influential variables $X_{13}$ and $X_{14}$ are independent $N(0,1)$, $(X_{15},X_{16})$ are jointly normal with mean zero, unit variances and covariances equal to $0.85$, and $X_{17},X_{18}$ and $X_{19}$ are independent variables with a $t$-distribution and 5 degrees of freedom.
+The features $X_{1}$ and $X_{6}$ are jointly normal with mean zero, unit variances, and covariance $0.85$.
+The features $X_{2},\ldots,X_{5}$ and $X_{7}$ are independent $N(0,1)$ variables, and $X_{8}$ is uniformly distributed on [1,1][-1,1].
+The binary feature $X_{9}=1\{X_{2}+Z<0\}$ is generated from an independent $Z\sim N(0,1)$.
+Feature $X_{10}$ follows a Poisson distribution with mean 3, while $X_{11}$ and $X_{12}$ are independent $t$-distributed variables with five degrees of freedom.
+The non-influential variables $X_{13}$ and $X_{14}$ are independent $N(0,1)$, $(X_{15},X_{16})$ are jointly normal with mean zero, unit variances, and covariance $0.85$, and $X_{17},X_{18}$, and $X_{19}$ are independent $t$-distributed variables with five degrees of freedom.
 We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples.
-We train a single-layer neural network model $f$ of the conditional mean $\mu(X)=\mathbb{E}(Y|X)$.
-Specifically,
+We train a single-layer neural network model $f$ to estimate the conditional mean $\mu(X)=\mathbb{E}(Y|X)$, defined as
 $f(x)=b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x),$ (14)
-where $K$ is the number of units, $\psi(y)=1/(1+e^{-y})$ is the activation function, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters to be estimated.
-The training procedure uses a standard $L_{2}$ loss criterion, $L_{2}$ regularization (of the weights $b_{k}$), the Adam optimization method, a batch size of 32, and 500 epochs with early stopping.
+where $K$ is the number of hidden units, $\psi(y)=1/(1+e^{-y})$ is the activation function, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters to be estimated.
+Training uses a standard $L_{2}$ loss criterion with $L_{2}$ regularization on the weights $b_{k}$, the Adam optimization method, a batch size of 32, and 500 epochs with early stopping.
 We iterate across five random initializations.
-A validation set ($25\%$ of $I_{1}$, not used for training) and a Bayesian approach are used to optimize the regularization weight, the learning rate and the number of units, $K$.
-The optimal weight is $7\cdot 10^{-4}$, the optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$.
-The $R^{2}$ on the training (and test) data exceeds $99\%$.
-The testing procedure uses a score function $S(m,y)=-(m-y)^{2}$ consistent with the $L_{2}$ training criterion.
-We use the basic baseline and comparison feature specifications discussed in Example 2.1.
-That is, for testing feature $\ell\in[d]$ we take
-$\underline{X}_{i}$ $=(X^{1}_{i},\ldots,X^{\ell-1}_{i},v^{\ell}_{i},X^{\ell+1}_{i},\ldots,X^{d}_{i})$
-$\overline{X}_{i}$ $=X_{i}=(X_{i}^{1},\ldots,X^{d}_{i})$ (15)
-where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ in case feature $\ell$ is continuous.
-For the one-hot feature $\ell=9$, we use $v^{\ell}_{i}=1-X^{\ell}_{i}$ and for the Poisson feature $\ell=10$, we use the “alternative mode” $v^{\ell}_{i}={\arg\max}_{x\in\mathbb{N}\setminus\{X^{\ell}_{i}\}}p^{\ell}(x|I_{1})$ where $p^{\ell}(\cdot|I_{1})$ is the probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$.
-The latter choice avoids baseline and comparison features being identical for too many samples, which is an issue arising with the standard mode.
-Our specifications yield baseline values in the support of $X^{9}$ and $X^{10}$.
+A validation set (25% of $I_{1}$, excluded from training) and a Bayesian search are used to optimize the regularization weight, learning rate, and number of units $K$.
+The optimal regularization weight is $7\cdot 10^{-4}$, the optimal learning rate is $5\cdot 10^{-4}$, and the optimal number of units is $K=300$.
+The resulting model achieves an $R^{2}$ exceeding 99% on both the training and test data.
+The testing procedure employs the score function $S(m,y)=-(m-y)^{2}$, consistent with the $L_{2}$ training criterion.
+We use the basic baseline and comparison feature specifications described in Example 2.1.
+For testing a feature $\ell\in[d]$, we set
+$\underline{X}_{i}$ $=(X^{1}_{i},\ldots,X^{\ell-1}_{i},v^{\ell}_{i},X^{\ell+1}_{i},\ldots,X^{d}_{i}),$
+$\overline{X}_{i}$ $=X_{i}=(X_{i}^{1},\ldots,X_{i}^{d}),$ (15)
+where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ when feature $\ell$ is continuous.
+For the one-hot feature $\ell=9$, we use $v^{\ell}_{i}=1-X^{\ell}_{i}$, and for the Poisson feature $\ell=10$, we adopt the “alternative mode”
+$v^{\ell}_{i}={\arg\max}_{x\in\mathbb{N}\setminus\{X^{\ell}_{i}\}}p^{\ell}(x|I_{1}),$
+where $p^{\ell}(\cdot|I_{1})$ denotes the probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$.
+This choice prevents baseline and comparison features from being identical for too many samples, an issue arising with the standard mode.
+Our specifications ensure that all baseline values lie within the support of $X^{9}$ and $X^{10}$.
 $\ell$ AICO CPT HRT LOCO Deep Gaussian
@@
 The test set size is $N=5\cdot 10^{5}$.
-We run 10 independent training/testing experiments to assess the behavior of the test.
-The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.999Limited numerical precision can lead to ties, i.e.
-samples $\Delta_{i}$ whose true values are distinct are being rounded to the same value.
+We run ten independent training/testing experiments to assess the behavior of the test.
+The second and third columns of Table 1 report results for confidence levels $\alpha=0.01$ and 0.050.05.777Limited numerical precision can lead to ties, i.e., samples $\Delta_{i}$ whose true values are distinct being rounded to the same value.
 If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$.
-Another numerical artifact is samples $\Delta_{i}$ with true, non-zero value being rounded to 0.
-We treat these samples as if they indeed had non-positive values not contributing to $n_{+}$, implying a conservative approach.
+Another numerical artifact arises when samples $\Delta_{i}$ with true, nonzero values are rounded to zero.
+We treat these as nonpositive, excluding them from $n_{+}$—a conservative choice.
 Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them.
@@
 for another approach.
-The test has perfect power and does not produce false discoveries.
-Feature correlation (e.g.
-$\ell=1,6$) and discrete feature distributions ($\ell=9,10$) do not affect performance.
-The test also correctly detects the significance of features $\ell=2,4,5$ which influence the response only through product terms.
-Additional experiments show robustness to alternative baseline specifications (e.g.
-taking $v^{\ell}_{i}$ as the training mean also for the discrete features $\ell=9,10$, and using the median instead of the mean to construct $v^{\ell}_{i}$).
-We also verify robustness with respect to the score function $S$ including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
+Across all experiments, the test achieves perfect power and produces no false discoveries.
+Feature correlation (e.g., $\ell=1,6$) and discrete feature distributions (e.g., $\ell=9,10$) do not impair performance.
+The procedure also correctly identifies the significance of features $\ell=2,4,5$, which influence the response only through interaction terms.
+Additional experiments confirm robustness to alternative baseline specifications—for example, taking $v_{i}^{\ell}$ as the training mean for the discrete features $\ell=9,10$, or using the median instead of the mean to construct $v_{i}^{\ell}$.
+We further verify robustness with respect to the score function $S$, including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
 […]
 […]
-Table 1 reports the results of the aforementioned alternative procedures for 10 independent training/testing experiments.
-CPT, HRT and LOCO deliver the same power as AICO but come with an increased false discovery rate (some but not all of these false discoveries can be addressed by a standard Bonferroni correction).
-The Gaussian Knockoff procedure selects the significant features 1 through 12 in each independent experiment.
-However, it does generate some false discoveries (which is consistent with control of the expected false discovery rate).
-The Deep Knockoff procedure reduces false discoveries at the expense of lower power.
-This approach has slight difficulty identifying the features entering the model (5.1) through the product term $2x_{4}x_{5}$.
-We consider the performance of our AICO test as a function of the size $N=|I_{2}|$ of the test set.
-Table 2 reports the results for $N\in\{10^{5},10^{4},10^{3},500,100\}$; the case $N=5\cdot 10^{5}$ is reported in Table 1.
-The test offers maximum power even for $N$ as small as 500.
-The power deteriorates moderately for smaller $N$; false discoveries rise only very slowly.
-This validates the performance of the test even for small samples.
+For additional perspective, we also implement strong model-agnostic variable-selection procedures, including the Gaussian Model-X Knockoff method of ?
+and the Deep Model-X Knockoff method of ?.
+The Gaussian version assumes a multivariate normal feature distribution, whereas the deep version uses a neural network to learn a generative model of the feature distribution.
+As with CPT and HRT, the construction of such auxiliary models incurs significant additional computational expense, which AICO avoids.
+Table 1 reports results for these alternative procedures across ten independent training/testing experiments.
+CPT, HRT, and LOCO achieve comparable power to AICO but exhibit higher false discovery rates (some of which, but not all, can be mitigated by a Bonferroni correction).
+The Gaussian Knockoff procedure consistently selects the significant features 1 through 12 in every experiment, though it produces a small number of false discoveries, consistent with its expected false-discovery-rate control.
+The Deep Knockoff method reduces false discoveries at the cost of lower power and shows modest difficulty identifying features that enter the model (5.1) through the interaction term $2x_{4}x_{5}$.
+We examine the performance of the AICO test as a function of the test-set size $N=|I_{2}|$.
+Table 2 reports results for $N\in\{10^{5},10^{4},10^{3},500,100\}$; the case $N=5\cdot 10^{5}$ is shown in Table 1.
+The test maintains full power for sample sizes as small as $N=500$.
+For smaller $N$, power declines moderately, while false discoveries increase only marginally.
+These results confirm that the test performs reliably even in small-sample settings.
 AICO
@@
 We report the number of times the null was rejected out of 10 trials for each covariate $\ell\in[d]$.
-We turn to measuring feature importance.
-Table 3 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen.
-It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities.
-Table 3 also provides the distribution of the randomized $p$-value.
-[^11] Ties of samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing CIs and coverage probabilities.
-This is consistent with their treatment in the testing procedure (footnote 9).
+We now turn to measuring feature importance.
+Table 3 presents a feature-importance ranking based on the empirical median (6) for one randomly selected training/testing experiment out of ten.
+It also reports the endpoints of the randomized confidence interval (7) for the population median $M$ at the $99\%$ level, along with the standard (non-randomized) two-sided interval (9) and their corresponding coverage probabilities.
+In addition, Table 3 provides the distribution of the randomized $p$-values.
+[^9] Ties among samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing confidence intervals and coverage probabilities.
+This treatment is consistent with the handling of ties in the testing procedure (footnote 7).
 AICO, Level $\alpha=1\%$
@@
 The table also reports the endpoints of the randomized CI (7) for the median at a $99\%$ level, the standard (non-randomized) two-sided CI (9) along with coverage probabilities, and the distribution of the randomized $p$-value.
-The $p$-value distributions show that our approach is able to clearly distinguish between the significant and irrelevant features.
-The distributions of the significant features $\ell=1,\ldots,12$ are concentrated near 0, far below the level $\alpha=1\%$.
-This provides substantial evidence in favor of rejecting the null for these features (the rejection probability is equal to 1).
-The $p$-value distributions of the irrelevant features are concentrated significantly above $\alpha$, implying clear retention of the null for these features.
-There are no equivocal cases, eliminating the need to actually randomize.
-Feature $\ell=9$ is judged to be the most important feature, followed by features $\ell=6,10$ and $12$.
-The remaining significant features are relatively less important compared with this group of variables.
-Features $\ell=4,5$ have (almost) identical medians and CIs and thus importance, consistent with their symmetric roles in the conditional mean function (5.1).
-To demonstrate the stability of AICO feature importance scores, Figure 4 shows the empirical median (6) for each of the ten training/testing experiments.
-The AICO importance ranking is stable and robust to noise in the data.
-The ranking is also robust to alternative choices of the score function $S$ such as S(m,y)=|my|S(m,y)=-|m-y| (results available upon request).
-For comparison, Figure 4 also provides test-implied feature importance scores for CPT, HRT, and LOCO (?, ?, ?), kernel and deep SHAP feature importance (?), permutation feature importance scores (?, ?), and Gaussian and deep knockoff feature importance scores (?, ?).
-It is important to note that there is no ground truth for feature importance (?); hence we do not focus on the differences in importance rankings across alternative approaches.
-We do, however, highlight the robustness of the measures based on testing and variable selection procedures.
-The widely-used SHAP measures deliver inconsistent rankings across implementations (deep vs.
-kernel).
-The rankings are also highly unstable across experiments (data sets).
-That is, they appear to be highly susceptible to noise in the data.
-This questions the usefulness of SHAP for feature importance.
+The $p$-value distributions show that our approach clearly separates significant from irrelevant features.
+For the significant features $\ell=1,\ldots,12$, the distributions are concentrated near zero, far below the level $\alpha=1\%$.
+This provides strong evidence in favor of rejecting the null hypothesis for these features, with rejection probability equal to one.
+In contrast, the $p$-value distributions of the irrelevant features are concentrated well above $\alpha$, indicating clear retention of the null.
+No equivocal cases are observed, eliminating the need for randomization.
+Feature $\ell=9$ emerges as the most important variable, followed by features $\ell=6$, $10$, and $12$.
+The remaining significant features are relatively less influential compared with this group.
+Features $\ell=4$ and $\ell=5$ exhibit (almost) identical medians and confidence intervals, consistent with their symmetric roles in the conditional mean function (5.1).
+To demonstrate the stability of AICO feature-importance scores, Figure 4 reports the empirical median (6) for each of ten independent training/testing experiments.
+The resulting AICO rankings are stable and robust to noise in the data, and remain consistent under alternative choices of the score function $S$, such as S(m,y)=|my|S(m,y)=-|m-y| (results available upon request).
+For comparison, Figure 4 also displays test-implied feature-importance scores for CPT, HRT, and LOCO (?, ?, ?); kernel and deep SHAP (?); permutation feature-importance scores (?, ?); and Gaussian and deep knockoff importance scores (?, ?).
+There is, of course, no ground truth for feature importance (?); we therefore focus on the stability of the rankings rather than their absolute differences across methods.
+The measures based on formal testing and variable-selection procedures, including AICO, exhibit strong robustness across experiments.
+In contrast, SHAP delivers inconsistent rankings between its deep and kernel implementations and shows substantial instability across data sets, indicating high sensitivity to noise.
+These findings raise concerns about the reliability of SHAP-based feature-importance measures.
 Figure 4: Feature importance scores across training/testing experiments.
-Finally we consider the computational efficiency of our test vs.
-the aforementioned alternatives.
-Figure 5 reports average computation times across 10 experiments.
-[^12] The compute environment consists of 8GB memory and 4 CPU cores (Intel Xeon CPU E5-2670, 2.60GHz and Silver 4110 CPU, 2.10GHz).
-In each of these experiments we test (AICO, CPT, HRT, LOCO) or screen (knockoffs) all 19 features or, in the case of SHAP and permutation feature importance, compute importance scores for all 19 features.
-The time reported includes the time to train any sub or auxiliary models.
-It does not include the time to train the original neural network (14).
+Finally, we compare the computational efficiency of our test with that of the alternative methods discussed above.
+Figure 5 reports average computation times across ten experiments.
+[^10] The compute environment consists of 8GB memory and four CPU cores (Intel Xeon CPU E5-2670, 2.60GHz, and Silver 4110 CPU, 2.10GHz).
+In each experiment we test (AICO, CPT, HRT, LOCO) or screen (knockoffs) all 19 features, or, in the case of SHAP and permutation feature importance, compute importance scores for all 19 features.
+The reported time includes any training of submodels or auxiliary models, but excludes the time required to train the original neural network (14).
 AICO achieves an unprecedented degree of computational efficiency.
-In terms of full-fledged testing procedures, AICO is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure.
-AICO is almost two orders of magnitudes more efficient than Gaussian knockoffs, the fastest screening procedure.
-It is more than three orders of magnitude faster than permutation importance, the fastest algorithmic feature importance measure.
-The testing results reported earlier in this section demonstrate that AICOs computational efficiency does not come at the expense of statistical performance.
+Among full-fledged testing procedures, it is more than three orders of magnitude faster than LOCO, the next most efficient test.
+It is nearly two orders of magnitude faster than Gaussian knockoffs, the fastest screening procedure, and more than three orders of magnitude faster than permutation importance, the most efficient algorithmic feature-importance method.
+As shown earlier in this section, AICO’s exceptional computational efficiency does not come at the expense of statistical performance.
 Figure 5: Average computation times for testing/screening all 19 features.
@@
 We next consider a binary response $Y\in\{0,1\}$ governed by the model
-$\mathbb{E}(Y|X)=\mathbb{P}(Y=1|X)=L(\mu(X))$ (16)
-where $L(u)=1/(1+e^{u})$, the function $\mu$ is given by (5.1) and the features $X=(X^{1},\ldots,X^{d})$ are as defined in Section 5.1.
+$\mathbb{E}(Y|X)=\mathbb{P}(Y=1|X)=L(\mu(X)),$ (16)
+where $L(u)=1/(1+e^{u})$, the function $\mu$ is given by (5.1), and the features $X=(X^{1},\ldots,X^{d})$ are defined as in Section 5.1.
 We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples.
-The data are imbalanced, as is often the case in practice (e.g.
-credit scoring): about $10\%$ of the samples have $Y_{i}=1$.
-We train a single-layer neural network model $f$ of the conditional mean $\mathbb{E}(Y|X)$:
-$f(x)=L\left(b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x)\right)$ (17)
-where $\psi(y)=1/(1+e^{-y})$, $K$ is the number of units, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters.
-The training procedure uses a standard cross-entropy criterion, $L_{2}$ regularization with weight $10^{-3}$ (approximately the optimal weight in the regression model training), Adam, a batch size of 32, and 500 epochs with early stopping.
-A validation set ($25\%$ of $I_{1}$) and a Bayesian approach are used to optimize the learning rate and the number of units, $K$.
+The data are imbalanced, as is common in practice (e.g., credit scoring): approximately 10% of the samples satisfy $Y_{i}=1$.
+We train a single-layer neural network $f$ to estimate the conditional mean $\mathbb{E}(Y|X)$, specified as
+$f(x)=L\!\left(b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x)\right),$ (17)
+where $\psi(y)=1/(1+e^{-y})$, $K$ is the number of hidden units, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters.
+Training uses a standard cross-entropy loss with $L_{2}$ regularization (weight $10^{-3}$, approximately optimal in the regression setting), the Adam optimizer, a batch size of 32, and 500 epochs with early stopping.
+A validation set (25% of $I_{1}$) and a Bayesian search procedure are used to tune the learning rate and number of units $K$.
 The optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$.
-The Area Under the ROC Curve on the training (and test) data exceeds $99\%$.
-The testing procedure employs a score function $S(m,y)=y\log m+(1-y)\log(1-m)$ consistent with the training criterion.
+The resulting model achieves an Area Under the ROC Curve exceeding 99% on both the training and test data.
+The testing procedure employs the score function $S(m,y)=y\log m+(1-y)\log(1-m)$, consistent with the training criterion.
 We use the baseline and comparison feature specification (5.1).
-Table 4 reports the results for confidence levels $\alpha=0.01,0.05$.
-As in the regression case, the test has perfect power and avoids false discoveries.
-The test is generally robust to alternative specifications of the baseline (e.g.
-using the median instead of the mean to construct $v^{\ell}$, setting $v^{\ell}$ to the training mean also for the 1-hot feature $\ell=9$).
-[^13] The local baseline has perfect power but some false discoveries.
-The test is also robust to using an alternative score function $S(m,y)=-(m-y)^{2}$.
+Table 4 reports the results for confidence levels $\alpha=0.01$ and $0.05$.
+As in the regression case, the test achieves perfect power and produces no false discoveries.
+It remains robust to alternative baseline specifications—for example, using the median instead of the mean to construct $v^{\ell}$, or setting $v^{\ell}$ to the training mean even for the one-hot feature $\ell=9$.
+[^11] The local baseline attains perfect power but yields some false discoveries.
+The test is also robust to alternative score functions, such as $S(m,y)=-(m-y)^{2}$.
 $\ell$ AICO CPT HRT LOCO Deep Gaussian
@@
 The test set size is $N=5\cdot 10^{5}$.
-Table 4 also reports the results for LOCO, CPT, HRT and Gaussian and deep knockoff procedures.
-AICO, CPT and LOCO have full power.
-LOCO, however, suffers from significant number of false discoveries (which cannot be addressed with a standard Bonferroni adjustment).
-HRT and knockoff procedures have less power and some false discoveries.
-In this comparison, AICO outperforms the alternatives.
-It performs as well as in the regression setting.
+Table 4 also reports results for LOCO, CPT, HRT, and the Gaussian and deep knockoff procedures.
+AICO, CPT, and LOCO all achieve full power.
+LOCO, however, produces a substantial number of false discoveries that cannot be corrected by a standard Bonferroni adjustment.
+HRT and the knockoff approaches exhibit lower power and some false discoveries.
+Overall, AICO outperforms all alternatives, matching its performance in the regression setting.
 AICO
@@
 We report the total number of rejections over 10 trials.
-Table 5 reports test results for varying test set sample sizes $N$.
-Power decreases faster with $N$ than in the regression case, especially for features $\ell=7,8$.
-False discoveries appear to be fairly insensitive to $N$.
+Table 5 reports test results for varying test-set sample sizes $N$.
+Power declines more rapidly with smaller $N$ than in the regression case, particularly for features $\ell=7$ and $\ell=8$.
+The incidence of false discoveries appears largely insensitive to $N$.
 AICO, Level $\alpha=1\%$
@@
 The table also reports the endpoints of the randomized CI (7) for the median at a $99\%$ level, the standard (non-randomized) two-sided CI (9) along with coverage probabilities, and the distribution of the randomized $p$-value.
-Table 6 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen.
-It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities.
-Table 3 also provides the distribution of the randomized $p$-value.
-Figure 6 shows the empirical median (6) for each of the ten training/testing experiments, for each of several methods including AICO.
-Across methods, the scores of the less important variables tend to be (somewhat) more volatile than in the regression setting.
-SHAP delivers inconsistent and highly unstable rankings, as in the regression case.
+Table 6 presents a feature-importance ranking based on the empirical median (6) for one randomly chosen training/testing experiment out of ten.
+It also reports the endpoints of the randomized confidence interval (7) for the population median $M$ at the $99\%$ level, together with the standard (non-randomized) two-sided interval (9) and their coverage probabilities.
+Table 3 further shows the distribution of the randomized $p$-values.
+Figure 6 displays the empirical median (6) across the ten independent training/testing experiments for several methods, including AICO.
+Across methods, the scores of the less important variables tend to be somewhat more volatile than in the regression setting.
+As before, SHAP yields inconsistent and highly unstable rankings across experiments.
 Figure 6: Classification model feature importance across training/testing experiments.
-Figure 7 reports average computation times across 10 experiments.
-As in the regression case, AICO achieves an unprecedented degree of computational efficiency.
+Figure 7 reports average computation times across ten experiments.
+As in the regression case, AICO achieves exceptional computational efficiency.
 It is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure.
-AICO is an order of magnitude more efficient than Gaussian knockoffs, the fastest screening procedure.
-It is more than three orders of magnitude faster than permutation importance, the fastest algorithmic feature importance measure.
+AICO is roughly an order of magnitude faster than Gaussian knockoffs, the fastest screening approach, and more than three orders of magnitude faster than permutation importance, the most efficient algorithmic feature-importance method.
 Figure 7: Average computation times for testing/screening all 19 features.
@@
 6.1 Credit card default
-This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.141414The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default.
-There are 23 explanatory variables (see Table 7).
-The dataset was split into training (60%), validation (15%), and testing (25%) sets with class proportions maintained through stratified sampling.
-Categorical variables are one-hot encoded and continuous variables are standardized (mean 0 and unit variance) based on training set statistics.
+This section complements the theoretical and numerical results with an empirical application to credit-card payment data from 30,000 card holders of a Taiwanese issuer, collected in October 2005.121212The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients.
+The response variable is a binary indicator of payment default, and there are 23 explanatory variables (see Table 7).
+The dataset is split into training (60%), validation (15%), and testing (25%) subsets, maintaining class proportions through stratified sampling.
+Categorical variables are one-hot encoded, and continuous variables are standardized to mean 0 and unit variance using training-set statistics.
 Feature Data Type Levels (for Categorical) Description
@@
 Table 7: Raw features of the credit default dataset.
-We implement an XGBoost (?) classifier using a cross-entropy loss function.
-We use the area under the ROC curve (AUC) on the validation set as a metric for hyper-parameter optimization.
-We employ the Optuna package (?), utilizing the Tree-Structured Parzen Estimator (?), with 250 trials.
-The search space includes the boosting algorithm, tree-specific settings such as tree depth and minimum child weight, and regularization terms.
-Class imbalance was mitigated using a scaling weight parameter, and training was capped at 1,000 boosting rounds with early stopping if the validation AUC did not improve for 50 consecutive rounds.
-The optimal hyper-parameters include a gradient tree boosting algorithm (?) with an approximation-based tree-building strategy; an L2 regularization parameter (lambda) of 0.2 and an L1 regularization parameter (alpha) of 0.001; a subsampling ratio of 100% (no subsampling); a column sampling ratio for each tree of 65%; a maximum tree depth of six levels; a minimum child weight of two; a learning rate (eta) of 0.05; a minimum loss reduction (gamma) of 0.00004; and a loss-guided tree growth policy.
-We train an ensemble of 100 XGBoost models on 100 bootstrap samples of the training data, all using the same hyper-parameters obtained during the tuning phase.
-Each model was trained for up to 1,000 boosting rounds, with early stopping triggered after 50 rounds of no improvement in the validation AUC.
+We implement an XGBoost classifier (?) using a cross-entropy loss function.
+Model performance is evaluated by the area under the ROC curve (AUC) on the validation set, which serves as the criterion for hyper-parameter optimization.
+We use the Optuna package (?) with the Tree-Structured Parzen Estimator (?), conducting 250 trials.
+The search space includes the boosting algorithm, tree-specific parameters such as maximum depth and minimum child weight, and regularization terms.
+Class imbalance is mitigated using a scaling-weight parameter, and training is capped at 1,000 boosting rounds with early stopping after 50 rounds without improvement in validation AUC.
+The optimal configuration employs gradient tree boosting (?) with an approximation-based tree-building strategy, an $L_{2}$ regularization parameter (lambda) of 0.2, an $L_{1}$ regularization parameter (alpha) of 0.001, a subsampling ratio of 100%, a column-sampling ratio per tree of 65%, a maximum tree depth of six, a minimum child weight of two, a learning rate (eta) of 0.05, a minimum loss reduction (gamma) of 0.00004, and a loss-guided tree-growth policy.
+We train an ensemble of 100 XGBoost models on 100 bootstrap samples of the training data, all using the tuned hyper-parameters.
+Each model is trained for up to 1,000 boosting rounds, with early stopping triggered after 50 rounds of no improvement in the validation AUC.
 Ensemble predictions are obtained by averaging the outputs of all 100 models.
-The resulting ensemble model has AUCs of $0.837$ on the training set, $0.782$ on the validation set, and $0.786$ on the test set.
-(The results are robust with respect to the model.
-An ensemble of neural networks, for example, yields very similar results.)
-As in the classification example of Section 5.2, we use a score function $S(m,y)=y\log m+(1-y)\log(1-m)$ to run the test.
-We continue to use the baseline and comparison feature specification (5.1), where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ in case feature $\ell$ is continuous.
-For a categorical variable we take the alternative mode $v_{i}^{\ell}=\arg\max_{x\in\mathcal{X}^{\ell}\setminus\{X_{i}^{\ell}\}}p^{\ell}(x|I_{1})$ where $\mathcal{X}^{\ell}$ is the (marginal) support of $X^{\ell}$ and $p^{\ell}(\cdot|I_{1})$ is the (empirical) probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$.
-This specification yields values in the support of a categorical.
-Here, we treat a categorical variable as a single feature rather than treating the associated one-hot variables separately.
-Alternatively, if higher resolution is required, we can run the test for each individual level of a categorical.
-Table 8 summarizes the test results for $\alpha=1\%$ ($N=7,500$ throughout).
-The significant features, ranked according to the empirical median (6), are PAY_0 through PAY_6, MARRIAGE, SEX, and AGE.
-Past delinquencies are significant predictors of future payment default.
-The more recent the payment record the stronger is the influence of timely payment.
-The actual monthly bill and payment amounts (e.g., PAY_AMT and BILL_AMT) are not significant.
-Marital status, gender and age, but not education level, are significant predictors of default, even after controlling for past payment behavior.
-However, the influence of these variables is relatively small, especially for age (the associated medians are at least an order of magnitude smaller than the medians of the payment records).
-These results indicate that personal circumstances have a statistically significant influence on borrower behavior, beyond patterns in prior payment.
+The resulting ensemble achieves AUCs of $0.837$ on the training set, $0.782$ on the validation set, and $0.786$ on the test set.
+The results are robust to the model choice; for example, an ensemble of neural networks yields nearly identical outcomes.
+As in the classification example of Section 5.2, we use the score function $S(m,y)=y\log m+(1-y)\log(1-m)$ to conduct the test.
+We continue to apply the baseline and comparison feature specification (5.1), where $v_{i}^{\ell}$ is the mean of $\{X_{k}^{\ell}:k\in I_{1}\}$ when feature $\ell$ is continuous.
+For categorical variables, we use the alternative mode
+$v_{i}^{\ell}=\arg\max_{x\in\mathcal{X}^{\ell}\setminus\{X_{i}^{\ell}\}}p^{\ell}(x|I_{1}),$
+where $\mathcal{X}^{\ell}$ denotes the marginal support of $X^{\ell}$ and $p^{\ell}(\cdot|I_{1})$ is the empirical probability mass function of $\{X_{k}^{\ell}:k\in I_{1}\}$.
+This specification ensures that the baseline value lies within the categorical support.
+We treat each categorical variable as a single feature rather than analyzing its associated one-hot encodings separately.
+If higher resolution is desired, the test can alternatively be run for each individual level of a categorical variable.
+Table 8 summarizes the test results for $\alpha=1\%$ (N=7,500N=7{,}500 throughout).
+The significant features, ranked by the empirical median (6), are PAY_0 through PAY_6, MARRIAGE, SEX, and AGE.
+Past delinquencies are strong predictors of future payment default, with more recent payment records exerting greater influence.
+Actual monthly bill and payment amounts (e.g., PAY_AMT and BILL_AMT) are not significant.
+Marital status, gender, and age—but not education level—are significant predictors of default even after controlling for past payment behavior.
+However, their influence is relatively small, especially for age: the corresponding medians are at least an order of magnitude smaller than those for the payment-history features.
+These findings suggest that personal circumstances exert a statistically significant, but comparatively modest, influence on borrower behavior beyond patterns in prior payment.
 AICO, Level $\alpha=1\%$
@@
 Table 8: AICO test summary for the credit card default dataset.
-We can further explore the role of a feature by running the test on subsets of the test set.
-For example, Table 9 reports the results for a test on the 4,029 (out of 7,500) samples for which $\texttt{MARRIAGE}=\text{Single}$.
-While EDUCATION is not significant at the $1\%$ level, it is significant at the $5\%$ level (the distribution of the $p$-value is concentrated below $0.05$).
-Furthermore, EDUCATION is insignificant when running the test on the samples for which $\texttt{MARRIAGE}=\text{Married}$ (results available on request).
-This means educational level is a significant predictor of default for an unmarried borrower but not a married one.
+We can further explore the role of individual features by running the test on subsets of the test set.
+For instance, Table 9 reports results for the 4,029 (out of 7,500) samples for which $\texttt{MARRIAGE}=\text{Single}$.
+While EDUCATION is not significant at the $1\%$ level, it becomes significant at the $5\%$ level, as the distribution of its $p$-value is concentrated below $0.05$.
+When the test is restricted to samples with $\texttt{MARRIAGE}=\text{Married}$ (results available upon request), EDUCATION is no longer significant.
+These findings indicate that educational level is a significant predictor of default for unmarried borrowers but not for married ones.
 AICO, Conditional on MARRIAGE = Single, Level $\alpha=1\%$
@@
 Table 9: AICO test summary for the credit card default dataset. The test is conditioned the samples for which $\texttt{MARRIAGE}=\text{Single}$.
-In the US, marital status, gender and age are protected characteristics under Fair Lending laws, so these variables are not used (they might not even be collected) to construct default probability estimates or credit scores for purposes of credit underwriting.
-Our results indicate, within the context of Taiwanese borrowers, that these attributes represent signals useful for risk prediction.
-All else equal, a score constructed without this set of variables might be less accurate.
+In the United States, marital status, gender, and age are protected characteristics under Fair Lending laws and therefore are not used—indeed, often not even collected—for estimating default probabilities or credit scores in credit underwriting.
+Our findings suggest that, within the context of Taiwanese borrowers, these attributes carry predictive information relevant to risk assessment.
+All else equal, a score constructed without these variables may be less accurate.
 6.2 Mortgage risk
-We next consider AICO feature significance for a more complex classification model trained to predict the state of a mortgage loan in the next month given a history of a time-varying feature vector available at the beginning of the month.
-The 7 states are Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent, Foreclosure, Paid Off, and Real Estate Owned.
-The analysis is based on an extensive panel dataset of mortgages from 4 ZIP codes in the greater Los Angeles area which are tracked monthly between 1/1994 and 12/2022 (this is a subset of the commercial CoreLogic Loan-Level Market Analytics dataset).
-The feature vector is of dimension $d=52$, and includes variables such as the current state, FICO credit score, loan balance, current interest rate, market interest rate, unemployment rate, etc., many of which are time-varying.
-We use the data filtering procedure of ?, resulting in roughly 5 million loan-month transitions from 117,523 active loans.
-The train/validation/test splits are as follows: 1/1994 - 6/2009, 7/2009 - 12/2009, and 1/2010 - 12/2022.
-We adopt the set-sequence model of ?
-with LongConv (?) for the sequence layer, which represents the current state of the art for mortgage transition data.
-We closely follow the training, validation and testing approach of ?.
+We next examine AICO feature significance in a more complex classification setting, predicting the state of a mortgage loan in the next month given a history of time-varying features observed at the beginning of the month.
+The seven states are Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent, Foreclosure, Paid Off, and Real Estate Owned.
+The analysis is based on an extensive panel dataset of mortgages from four ZIP codes in the greater Los Angeles area, tracked monthly from January 1994 through December 2022.
+This dataset constitutes a subset of the CoreLogic Loan-Level Market Analytics data.
+The feature vector has dimension $d=52$ and includes variables such as current loan state, FICO credit score, loan balance, current interest rate, market interest rate, and local unemployment rate, many of which are time-varying.
+We apply the data-filtering procedure of ?, resulting in roughly five million loan-month transitions from 117,523 active loans.
+The training, validation, and test splits are 1/1994–6/2009, 7/2009–12/2009, and 1/2010–12/2022, respectively.
+We adopt the set–sequence model of ?
+with LongConv (?) for the sequence layer, which represents the current state of the art for mortgage transition modeling.
+Training, validation, and testing procedures follow closely the approach of ?.
 The dataset contains missing feature values, which are handled using missingness dummies and forward-filling based on the last observed value for each loan.
@@
 This paper develops a simple and computationally efficient framework for assessing feature significance in supervised learning.
-Our randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation.
-Empirical studies show strong performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, demonstrating its ability to inform high-stakes decisions with legal, economic, and regulatory implications.
-More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines statistical grounding with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
+The proposed randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation.
+Empirical studies demonstrate strong statistical performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, illustrating the method’s ability to inform high-stakes decisions with legal, economic, and regulatory implications.
+More broadly, the framework establishes a foundation for a new standard of interpretability in machine learning—one that combines statistical rigor with computational scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
 Appendix A Proofs of technical results
@@
 Appendix B Panel data
-Consider panel data $\{Z_{i,t}\}$ with $i=1,\ldots,m$ indexing units and $t=1,\ldots,T$ indexing time.
+Consider panel data $\{Z_{i,t}\}$, where $i=1,\ldots,m$ indexes units and $t=1,\ldots,T$ indexes time.
 We partition $\{(i,t):i\in[m],t\in[T]\}$ into two disjoint subsets, $I_{1}$ and $I_{2}$, such that each pair belongs to exactly one subset.
-This formulation allows for commonly used data-splitting schemes, such as temporal splits (distinguishing between in-sample and out-of-sample time periods), unit splits (assigning a unit’s entire time series to either $I_{1}$ or $I_{2}$), or combinations of both.
-In the case of an unbalanced panel, where units may be observed at different points in time, only observed pairs $(i,t)$ are assigned to $I_{1}$ or $I_{2}$.
+This setup accommodates common data-splitting schemes, including temporal splits (distinguishing between in-sample and out-of-sample time periods), unit-based splits (assigning each unit’s entire time series to either $I_{1}$ or $I_{2}$), or combinations of both.
+For an unbalanced panel, in which units may be observed at different times, only observed pairs $(i,t)$ are assigned to $I_{1}$ or $I_{2}$.
 B.1 Conditioning
-Conditioning entails focusing on an appropriate subset $I\subseteq I_{2}$ when constructing the feature effect.
-For example, in the credit card analysis of Section 6.1, we condition on card applicants who are married to understand feature significance for a group of borrowers with certain features.
-If some samples in $I_{2}$ are dependent, as is often the case for panel data, we can choose $I$ as the subset of independent samples $(i,t)\in I_{2}$ (albeit at the cost of reducing the size of the test set).
-We can further partition $I$ by period $t$, to construct feature effects $\{\Delta_{\cdot,t}:(\cdot,t)\in I\}$ indexed by $t$.
-This approach leads to a sequence of tests, and a time series of $p$-values as well as importance scores and their CIs.
-The time-series behavior may provide useful insight.
-For the purpose of constructing baseline and comparison features, the training set $I_{1}$ is also treated as a collection of $|I_{1}|$ samples–with unit and time information tied to each sample–fully harnessing panel data richness.
-For each $(i,t)\in I_{2}$, the baseline features can harness information from not only all training samples $I_{1}$ but also testing samples corresponding to the same unit $\{X_{i,\cdot}:(i,\cdot)\in I_{2}\}$.
-This flexibility enables unit-specific and time-dependent baselines such as lagged feature value (e.g., last month’s feature value) and historical aggregation, enhancing the baseline’s meaningfulness.
+Conditioning entails focusing on a suitable subset $I\subseteq I_{2}$ when constructing the feature effect.
+For example, in the credit card analysis of Section 6.1, we condition on card applicants who are married to assess feature significance within a specific borrower subgroup.
+If some samples in $I_{2}$ are dependent—as is common in panel data—we can restrict attention to the subset $I$ of independent samples $(i,t)\in I_{2}$, albeit at the cost of a smaller test set.
+We can further partition $I$ by period $t$ to construct feature effects $\{\Delta_{\cdot,t}:(\cdot,t)\in I\}$ indexed by time.
+This yields a sequence of tests and a time series of $p$-values, importance scores, and confidence intervals whose temporal dynamics may provide additional insight.
+For constructing baseline and comparison features, the training set $I_{1}$ is treated as a collection of $|I_{1}|$ samples—with unit and time information tied to each—fully exploiting the richness of the panel structure.
+For each $(i,t)\in I_{2}$, the baseline features can draw information not only from all training samples $I_{1}$ but also from testing samples corresponding to the same unit, $\{X_{i,\cdot}:(i,\cdot)\in I_{2}\}$.
+This flexibility enables unit-specific and time-dependent baselines, such as lagged feature values (e.g., the previous month’s feature value) or historical aggregations, thereby enhancing the contextual relevance of the baseline.
 B.2 Grouping
-An alternative grouping approach allows for joint testing across the time-series.
-The idea is to aggregate feature effects within a unit of dependence, so that dependence is contained within groups.
-We treat the testing set $I_{2}$ as a collection of $|I_{2}|$ samples and compute the baseline and treatment predicted responses: $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, $(i,t)\in I_{2}$, where $\underline{X}_{i,t}$ and $\overline{X}_{i,t}$ are baseline and treatment feature values constructed as in the conditional method.
-Note that this differs slightly from the conditional approach.
-In the grouping approach, $I_{2}$ is treated as a collection of $|I_{2}|$ samples only until the construction of the baseline and comparison predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, unlike the conditional approach which includes the construction of the feature effect $\Delta_{i,t}$.
-Then, we group the computed predictions along with the corresponding true responses by unit to compute the unit-level feature effect
-$\Delta_{i}=S(\overline{f}_{i},Y_{i})-S(\underline{f}_{i},Y_{i})$
+An alternative grouping approach enables joint testing across time series.
+The idea is to aggregate feature effects within a unit of dependence so that dependence is contained within groups.
+We treat the testing set $I_{2}$ as a collection of $|I_{2}|$ samples and compute the baseline and treatment predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$ for $(i,t)\in I_{2}$, where $\underline{X}_{i,t}$ and $\overline{X}_{i,t}$ are baseline and treatment feature values constructed as in the conditional method.
+This differs slightly from the conditioning approach: in the grouping approach, $I_{2}$ is treated as a collection of $|I_{2}|$ samples only up to the computation of the baseline and comparison predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, whereas the conditional approach proceeds to construct the feature effect $\Delta_{i,t}$ at the individual-sample level.
+Next, we group the computed predictions and corresponding true responses by unit to obtain the unit-level feature effect
+$\Delta_{i}=S(\overline{f}_{i},Y_{i})-S(\underline{f}_{i},Y_{i}),$
 where $S$ is a trajectory-level score function, $\underline{f}_{i}=\{f(\underline{X}_{i,t}):(i,t)\in I_{2}\}$, $\overline{f}_{i}=\{f(\overline{X}_{i,t}):(i,t)\in I_{2}\}$, and $Y_{i}=\{Y_{i,t}:(i,t)\in I_{2}\}$.
-Unlike previous single-period settings where the score function $S$ evaluates only one prediction against one true response, the trajectory-level score function evaluates the entire time series of predictions against the time series of true responses.
-The trajectory-level score function can extend single-period score functions.
-An example is the negative mean squared error, where the average is across time steps.
-More complex score functions can harness rich temporal data, like correlation (Information Coefficient, IC), negative error in maximum drawdown (MDD): S(m,y)=|MDD(y)MDD(m)|S(m,y)=-|MDD(y)-MDD(m)|, or other time series evaluations suitable for specific domain usage.
+Unlike the single-period setting, where $S$ evaluates one prediction against one true response, the trajectory-level score function evaluates the entire sequence of predictions against the corresponding sequence of true responses.
+Trajectory-level score functions can extend familiar single-period metrics.
+For instance, one may use the negative mean squared error, averaging across time steps.
+More complex specifications can exploit richer temporal structure, such as correlation (Information Coefficient, IC), the negative error in maximum drawdown (MDD), S(m,y)=|MDD(y)MDD(m)|S(m,y)=-|MDD(y)-MDD(m)|, or other time-series evaluations tailored to specific domains.
 More broadly, the grouping approach extends beyond temporal dependence structures.
-For instance, for spatially dependent data, one could group testing samples by region or neighborhood, under the assumption that neighborhoods are (approximately) independent.
-Similarly, in settings with repeated measurements, grouping by the underlying entity or context may be effective.
+For spatially dependent data, one may group testing samples by region or neighborhood, under the assumption that neighborhoods are approximately independent.
+Likewise, for repeated-measures data, grouping by the underlying entity or experimental context can provide an effective way to respect within-group dependence while maintaining valid inference across groups.
 B.3 Hybrid approach
 Conditioning and grouping can be combined.
-For example, suppose we want to analyze a feature’s impact on model performance over a period of time.
-A long period avoids short-term noise but renders single-period conditioning insufficient.
-A shorter period ensures the relevance of data but may rule out grouping.
-This can be addressed by first conditioning on a time frame of desirable length and then grouping the conditional data by units corresponding to each time-unit pair.
-As with conditioning, this procedure can be repeated across different time windows, leading to a rolling-window analysis with smoothed single-period results.
-Another notable use case of the hybrid method is the analysis of per-unit consistency of feature performance: i.e., for each unit, how consistent across time is a feature’s impact on model performance?
-A longer window reduces short-term fluctuations; when the multi-period score function is an aggregation of some single-period scores, a longer window also introduces another interesting aspect: per-unit consistency.
-For example, suppose we choose the multi-period score function to simply be the average of a single-period function, such as negative mean squared error.
-Then, it can be shown that the unit-level feature effect $\Delta_{i}$ is essentially the average of unit-time-level feature effects $\Delta_{i,t}$ and, therefore, embeds some degree of consistency across time.
-The longer the time window, the higher the degree of consistency imposed, as the bar is set higher.
+For instance, suppose we wish to analyze a feature’s impact on model performance over an extended period.
+A longer period helps smooth short-term noise but makes single-period conditioning insufficient.
+Conversely, a shorter period preserves data relevance but may preclude grouping.
+This limitation can be addressed by first conditioning on a time frame of suitable length and then grouping the conditioned data by units corresponding to each time–unit pair.
+As with conditioning alone, this procedure can be repeated across overlapping time windows, yielding a rolling-window analysis with smoothed single-period results.
+Another notable use case of the hybrid method is assessing per-unit consistency of feature performance—how consistent over time is a feature’s impact on model behavior for a given unit?
+A longer window mitigates short-term fluctuations, and when the multi-period score function aggregates single-period scores, it also captures an additional dimension of per-unit consistency.
+For example, consider a multi-period score function defined as the average of a single-period function, such as the negative mean squared error.
+In this case, the unit-level feature effect $\Delta_{i}$ is effectively the average of the unit–time-level effects $\Delta_{i,t}$ and thus reflects a degree of consistency over time.
+The longer the time window, the stronger the imposed consistency, as a higher bar is set for variation in feature impact.
 Appendix C Missing Data